### v1

# RoboTransfer: Geometry-Consistent Video Diffusion for Robotic Visual Policy Transfer

Abstract
Imitation Learning has become a fundamental approach in robotic manipulation. However, collecting large-scale real-world robot demonstrations is prohibitively expensive. Simulators offer a cost-effective alternative, but the sim-to-real gap make it extremely challenging to scale. Therefore, we introduce RoboTransfer, a diffusion-based video generation framework for robotic data synthesis. Unlike previous methods, RoboTransfer integrates multi-view geometry with explicit control over scene components, such as background and object attributes. By incorporating cross-view feature interactions and global depth/normal conditions, RoboTransfer ensures geometry consistency across views. This framework allows fine-grained control, including background edits and object swaps. Experiments demonstrate that RoboTransfer is capable of generating multi-view videos with enhanced geometric consistency and visual fidelity. In addition, policies trained on the data generated by RoboTransfer achieve a 33.3% relative improvement in the success rate in the Diff-Obj setting and a substantial 251% relative improvement in the more challenging Diff-All scenario. Explore more demos on our project page: https://horizonrobotics.github.io/robot_lab/robotransfer
**footnotetext: Equal contribution. footnotetext: Corresponding author.
1 Introduction
Imitation Learning (IL) has become a fundamental approach for visuomotor control in robotic manipulation [67, 68]. However, collecting large-scale real-world robot demonstrations is prohibitively expensive [6, 7]. Simulated environments offer a cost-effective alternative [54, 59, 19, 44], but the scarcity of assets and sim-to-real gap[51, 43] make it extremely challenging to scale.
Recently, diffusion models [26] have gained attention as a promising method to synthetically generate realistic and diverse data. To expand robotic datasets without the need for extensive real-world data collection, ROSIE [62] employs text-guided image generation models trained on real-world data. However, it performs image augmentation on a per-frame basis, which leads to a loss of temporal and spatial consistency. In contrast, Cosmos-Transfer [2] generates photorealistic data using a video diffusion model conditioned on segmentation and depth, which helps preserve geometric consistency.
Despite the rapid advancements in the field, two key challenges remain. First, robotic systems often rely on observations from a single viewpoint, which inherently limits their perception. To overcome this limitation, multi-view observation is a commonly used approach. However, generating multi-view consistent results remains a significant challenge for video generative models. Second, robot manipulation scenarios are typically complex and highly interactive. Controlling such scenarios precisely through textual input alone, as demonstrated in text-to-video frameworks [5, 70], presents substantial difficulties.
In this work, we introduce RoboTransfer, as is shown in Figure 1, a video generation framework designed for data synthesis in robotics. To the best of our knowledge, it is the first framework to integrate multi-view geometry while offering explicit control over background, object attributes. RoboTransfer incorporates cross-view feature interactions and global depth/normal conditions within the video diffusion architecture, ensuring 3D consistency across different camera views and over time. To enhance controllability, RoboTransfer decomposes each scene into distinct control conditions, enabling background edits and object swaps by conditioning on this decomposition.
[…]
Figure 1: Overview of RoboTransfer.
Our work makes the following contributions:
  1. 1.
    We propose RoboTransfer, a robot generation framework based on multi-view, geometry-aware video diffusion models that maintain fidelity across multiple reference conditions.
  2. 2.
    We introduce a dataset construction pipeline that decomposes real-world recording data into high-quality triplets, which incorporate global geometry and appearance control conditions.
  3. 3.
    Through carefully designed experiments and evaluations, we demonstrate that our approach generates multi-view videos with significantly improved geometric consistency and visual fidelity. Furthermore, policies trained with RoboTransfer-generated data exhibit better generalization to novel and unseen environments.
2 Related Work
2.1 Visual Generalizable Imitation Learning
Imitation Learning (IL) has emerged as a foundational approach for visuomotor control in robotic manipulation. Deep networks trained on collected demonstrations—e.g., DP [12], OpenVLA [31], Act [67], Aloha [68], E2E-VP [36], and RT2 [7]—map raw visual inputs directly to motor commands. To bolster policy robustness and generalization, researchers have explored novel architectures [12] and vision pretraining schemes [4, 13]. However, gathering large-scale, task-specific robot data remains prohibitively expensive. Consequently, many methods leverage auxiliary sources—human demonstration videos [20], high-fidelity simulators [54, 59, 19, 44], or operational logs from other platforms [46, 30]. Each of these domains exhibits its own biases and visual characteristics, introducing distributional shifts that can degrade IL policy performance when transferred to the target environment. In this work, we offer a complementary perspective on generalizable visual imitation learning by leveraging state-of-the-art video diffusion models. By generating target-conditioned training data, our approach significantly enhances the robustness of policy and enables seamless transfer to novel environments.
2.2 Data Augmentation for Policy Learning
Data augmentation is widely used to train robust machine learning models. Prior work has applied augmentation directly to real-world robot captures to improve the resilience of visual policies against environmental variations[41, 42, 15, 23, 21]. Image-inpainting methods further enhance appearance invariance by replacing or masking out scene elements in collected frames [62, 11, 10, 40]. Likewise, neural view synthesis techniques generate novel viewpoints from existing robot data, promoting view-invariant feature learning [53, 8]. Simulators can cheaply produce vast quantities of labeled frames [54, 59, 19, 44], but discrepancies in physics modeling, rendering fidelity, and scene composition often hinder direct sim-to-real transfer. Domain randomization is used to close the sim-to-real gap, thereby enabling policies trained in simulation to generalize more effectively to real-world environments. [35, 22, 34, 51, 1, 48, 25]. In contrast to these methods, we leverage a video diffusion model to generate photorealistic image manipulations. This approach preserves both multi-view and geometry consistency, producing richly detailed samples for robot policy training.
2.3 Generation Models in Embodied AI
With the rapid advancement of generative models such as HunyuanVideo [33], Wan [56], SVD [5], Cosmos [2], CogVideoX [61], OpenSora [70], there is a growing trend toward leveraging generated data to bridge the gap between synthetic environments and real-world physical applications [72]. In particular, domains like autonomous driving have seen increasing adoption of generative techniques, as demonstrated by [57, 66, 27, 58, 17, 18, 64, 45, 65]. Similarly, robotic scenarios—where real-world data collection is often prohibitively expensive and time-consuming—have greatly benefited from generative data. Models such as UniSim [60], UniPi [14], and RoboDreamer [71] synthesize future robot behaviors via text prompts, which are then translated into actionable commands using inverse dynamics models. However, purely video-based generation methods often suffer from spatial and temporal inconsistencies, leading to degraded performance in downstream action prediction and planning. To address this, TesserAct [69] proposes a multi-modal generation pipeline that simultaneously synthesizes RGB, depth, and normal videos, enabling the reconstruction of coherent 4D scenes. This framework ensures both spatial and temporal consistency in robotic environments and supports policy learning that significantly surpasses prior video-only world models. Moreover, data generalization remains a critical challenge in robot learning. Techniques like ReBot [16] and RoboEngine [63] adopt background inpainting to diversify environmental textures, while Cosmos-Transfer [2] utilizes video-to-video translation to enrich overall scene appearance. Despite their success, these approaches lack fine-grained control over foreground and background textures, often resulting in synthetic data distributions misaligned with real-world tasks, especially when training policy models. Additionally, their inability to ensure multi-view consistency limits their scalability and effectiveness in complex, robotic settings.
3 Methods
We introduce RoboTransfer, a framework designed for the controllable generation of multi-view videos to support the training of policy models. By providing explicit control over both scene geometry and appearance, RoboTransfer enables the synthesis of video data with precisely defined distributions. This fine-grained control allows for the generation of diverse training scenarios, which are critical for improving the generalization of policy models. In the remainder of this section, we first present the preliminaries of video diffusion models in Sec. 3.1. Then, we describe the framework of RoboTransfer in Sec. 3.2. Finally, the dataset construction pipeline is elaborated in Sec. 3.2.
3.1 Video Diffusion Model Preliminaries
Diffusion-based methods for controllable video generation typically model the process as a gradual refinement of a noise latent $\boldsymbol{\epsilon}$ into a clean video latent $\mathbf{x_{0}}$, under the guidance of spatially aligned conditions $y_{s}$ (e.g., depth map, segmentation map) and unstructured conditions $y_{u}$ (e.g., CLIP embedding). Popular approaches in this paradigm include Flow Matching [37], DDPM [26], and EDM [29]. In the case of EDM, the learning objective is formulated as:
$\mathcal{L}(D_{\theta},\sigma)=\mathbb{E}_{\mathbf{x_{0}},y_{s},y_{u}}\left[\left\|\mathbf{x_{0}}-D_{\theta}\left(\mathcal{E}(\mathbf{x_{0}}+\mathbf{n}),\tau_{u}(y_{u}),\tau_{s}(y_{s}),\sigma\right)\right\|_{2}^{2}\right].$ (1)
Here, $\mathbf{x_{0}}\sim p_{\text{data}}$ denotes a clean sample drawn from the dataset, and $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ represents Gaussian noise. $\mathcal{E}$ denotes the encoder component of a VAE [32], and the denoising network $D_{\theta}$ is conditioned on the noise level $\sigma$ and encoded conditions $\tau_{u}(y_{u}),\tau_{s}(y_{s})$. To encourage stable learning across varying noise magnitudes, the total training objective aggregates the per-noise loss via a weighted expectation:
$\mathcal{L}(D_{\theta})$ $=\mathbb{E}_{\sigma}\left[\frac{\lambda(\sigma)}{\exp(u(\sigma))}\mathcal{L}(D_{\theta},\sigma)+u(\sigma)\right],$ (2)
$\lambda(\sigma)$ $=\frac{\sigma^{2}+\sigma_{\text{data}}^{2}}{(\sigma\cdot\sigma_{\text{data}})^{2}},$ (3)
$\ln(\sigma)$ $\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2}).$ (4)
In this setup, $\sigma_{\text{data}}$ denotes the empirical standard deviation of the data, while the distribution of $\sigma$ is governed by the hyperparameters $P_{\text{mean}}$ and $P_{\text{std}}$. The weighting function $\lambda(\sigma)$ ensures that all noise levels contribute proportionally during the initial stages of training.
[…]
Figure 2: The RoboTransfer framework performs multi-view consistent modeling to jointly reason across viewpoints. It represents geometry with scaled depth and normal maps, and encodes appearance using reference backgrounds and object-specific images for detailed control over appearance.
3.2 RoboTransfer Framework
The overall framework of RoboTransfer is illustrated in Figure 2. To ensure multi-view consistency during generation, we perform multi-view consistent modeling, enabling the generation process to reason jointly over information from different viewpoints. On the conditions side, RoboTransfer incorporates fine-grained control by encoding both geometric and appearance information. Specifically, we represent geometry using scaled depth maps and surface normal maps, capturing the underlying 3D structure of the scene. Meanwhile, the appearance is encoded using reference background images and object-specific images, providing detailed control over texture, color, and contextual appearance. In the following sections, we first introduce the multi-view consistent modeling, and then describe the encoding mechanisms for geometric and appearance conditions, respectively.
Multi-view Consistent Modeling. In robotic environments, it is common to deploy multiple cameras from different viewpoints to capture a scene simultaneously. To generate multi-view consistent videos for downstream tasks, RoboTransfer leverages the multi-view in-context learning capabilities [66, 28] inherently present in pretrained diffusion models. Specifically, given $N$ synchronized videos $\{V_{1},V_{2},...,V_{N}\}$ from different viewpoints, we concatenate them along the width dimension and encode the joint information using a VAE encoder $\mathcal{E}$:
$\mathbf{x}_{0}=\mathcal{E}([V_{1},V_{2},\ldots,V_{N}]),$ (5)
where $[V_{1},V_{2},\ldots,V_{N}]$ represents the concatenated multi-view video sequence. This modeling approach maintains the integrity of the pretrained video diffusion architecture, requiring no structural modifications, while enabling fast convergence and high-quality, view-consistent video generation in robotic settings.
Geometry Conditions. Recent video generation models primarily learn spatiotemporal coherence through data-driven approaches such as temporal attention [5, 70] or full attention [56, 33, 2]. While effective at capturing pixel-level dynamics, these methods lack an explicit understanding of underlying 3D geometry, limiting their applicability in robotic environments. To address this limitation, RoboTransfer explicitly incorporates geometry conditions to enhance the model’s awareness of scene structure and depth continuity. Specifically, we utilize depth and surface normal videos as geometric conditions. These two types of geometric cues are complementary — depth maps provide information about spatial distances from the camera, while normal maps encode local surface orientations. When combined, they offer a richer and more holistic description of scene geometry (see Sec. 3.2 for details on data acquisition). Since both the depth and normal videos are spatially aligned with the RGB views, we propose GuideNet, which utilizes a stack of convolutional layers that jointly downsample and encode the depth and normal sequences. The resulting geometry-aware representation is then concatenated with the noise latents along the channel dimension, allowing the generation process to be guided by accurate 3D structural cues.
Appearance Conditions. To enable fine-grained control over the generated textures, RoboTransfer introduces texture conditions from two complementary perspectives: background appearance and object appearance. Specifically, we use a background reference image and a set of object reference images as conditions signals to guide the generation process toward faithful reproduction of scene textures and object-level details. A key challenge in appearance control is ensuring compatibility with the geometry conditions. Naïvely combining appearance and geometry inputs may introduce conflicts (e.g., mismatched depth and texture), thereby weakening both the structural consistency and the visual fidelity of generated results. To address this, we carefully curate the appearance reference images, resulting in a background reference image $C_{b}$ and object reference images $C_{o}$ that do not contradict the geometric priors (see Sec. 3.2 for details). For background appearance, we employ a VAE encoder to compress $C_{b}$ into a latent representation that is spatially aligned with the generation latents. This encoded background appearance is then concatenated with the latent inputs to control the global texture and background style of the output video. In contrast, object appearance presents additional challenges due to the variable number and spatial distribution of objects in the scene. Therefore, object images are treated as unstructured conditions. We encode each object’s appearance using CLIP [47], which produces a global embedding used to guide the generation via cross attention. This design allows flexible and scalable conditioning on diverse object appearances while preserving compatibility with the structured geometry inputs.
[…]
Figure 3: RoboTransfer data processing pipeline. The data processing pipeline consists of two main components: Geometry conditions (left) are derived by using Mono-Normal and Video Depth Anything, which are scale-aligned with sensor depth to ensure consistency. Appearance conditions (right) are obtained by sampling keyframes from the video. GPT-4 is used to generate object descriptions, which are then processed by Grounding-SAM to create per-object masks. Additionally, background inpainting is used to generate complete reference backgrounds.
3.3 Dataset Construction from Real World Data
Data construction is central to our framework. The overall pipeline is illustrated in Figure 3. Unlike conventional text-to-video or image-to-video tasks, our approach requires both globally consistent geometry and multiple part-specific reference images (e.g., background and manipulable objects).
Our pipeline begins with the collection of a large-scale robotic demonstration dataset. Each demonstration captures synchronized RGB-D streams from both a hand-eye and an overhead RGB-D camera. To cover a wide range of operating conditions, we vary backgrounds (e.g., textured tabletops, cluttered surfaces) and manipulable objects (shapes, sizes, materials).
To enforce spatial and temporal consistency, we incorporate geometric cues such as depth and surface normals. Since raw depth maps from RGB-D sensors are often noisy, we employ the state-of-the-art video depth estimator, Video Depth Anything[9], to produce temporally consistent depth maps. We then align the scale of the estimated depth map with the RGB-D sensor’s depth using robust least-squares fitting to ensure global spatial accuracy. Additional details are provided in the supplementary material. For surface normals, we compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS[24], which jointly serves as reliable geometric conditions.
Next, we sample a subset of RGB keyframes as visual references. These frames are processed by GPT-4 to generate structured scene descriptions, including object-level descriptions for individual entities. These descriptions guide the Grounding-SAM module[50], which detects and segments per-object masks. To obtain clean background references, we inpaint the segmented objects from the original images, generating plausible object-free scenes (e.g., an empty tabletop). For multimodal object-level conditions, the segmented object masks are further processed with the CLIP model[47] to obtain semantic embeddings. These conditions allow the diffusion model to generate photorealistic, semantically grounded video augmentations that respect both object identity and spatial layout.
4 Experiments
4.1 Synthesis Quality
Implementation Details. To construct our training dataset, we leverage the open-source Cobot Magic platform to collect a large-scale video corpus of dual-arm robot executions, covering diverse tasks such as grasping and placing. The raw videos are segmented into 10Hz clips of 30 frames each, resulting in approximately 24k clips for training. Additionally, we curate a set of 1.6k 10Hz 30-frame clips from the collected dataset for video synthesis quality evaluation. All videos are annotated with conditions as described in Sec. 3.2, to facilitate both training and evaluation. RoboTransfer is fine-tuned from the pretrained video diffusion model [5]. During training, videos from each camera view are resized to a resolution of $640\times 384$. We adopt the AdamW optimizer with a learning rate of $3\times 10^{-5}$ and a global batch size of 8, training for a total of 70K steps. During inference, we use the EDM scheduler [29] to perform 30 denoising steps and apply classifier-free guidance.
Table 1: Experiment comparison on different geometry conditions of RoboTransfer.
Model Camera RMSE $\downarrow$ Abs.Rel. $\downarrow$ Sq.Rel. $\downarrow$ Mean Err. $\downarrow$ Med.Err. $\downarrow$ Pix.Mat. $\uparrow$ FVD $\downarrow$
RoboTransfer [D.S.] left 0.074 0.124 0.020 4.86 2.88 142.90 218.51
RoboTransfer [D.P.] left 0.054 0.090 0.010 3.91 2.28 149.68 123.31
RoboTransfer [Scaled D.P.] left 0.049 0.081 0.008 3.48 1.99 183.26 112.43
RoboTransfer [Scaled D.P. + N.] left 0.047 0.079 0.008 3.31 1.92 202.03 107.43
RoboTransfer [D.S.] middle 0.182 0.074 0.031 3.51 1.58 153.76
RoboTransfer [D.P.] middle 0.132 0.053 0.015 3.05 1.43 95.89
RoboTransfer [Scaled D.P.] middle 0.134 0.054 0.015 2.93 1.40 103.32
RoboTransfer [Scaled D.P. + N.] middle 0.133 0.054 0.015 2.86 1.39 101.17
RoboTransfer [D.S.] right 0.090 0.137 0.025 4.93 2.91 40.70 396.33
RoboTransfer [D.P.] right 0.072 0.103 0.016 4.14 2.36 56.02 262.96
RoboTransfer [Scaled D.P.] right 0.064 0.090 0.011 3.74 2.11 65.45 226.76
RoboTransfer [Scaled D.P. + N.] right 0.058 0.087 0.009 3.55 2.00 75.67 220.12
Evaluation Metrics. We evaluate the generated videos from multiple perspectives, including multi-view consistency, geometric consistency, and semantic consistency. For multi-view consistency, we follow [3] and utilize the state-of-the-art image matching method [52] to compute the number of matched pixels (Mat.Pix.) between adjacent views (i.e., left-to-center and right-to-center). For geometric consistency, we adopt the evaluation framework from Cosmos-Transfer [2], which assesses depth and surface normal alignment. For depth prediction, we compute scale-invariant metrics including Root Mean Squared Error (RMSE), Absolute Relative Error (Abs.Rel.), and Squared Relative Error (Sq.Rel.). For normal estimation, we report the Mean Angular Error (Mean Err.) and Median Angular Error (Med. Err.). In terms of semantic consistency, we measure appearance controllability using CLIP [47] similarity. For background-level similarity, we compute the CLIP cosine similarity between a reference background image and each generated frame (BG. Sim.). For foreground object similarity, we use GroundingDINO [38] and SAM2 [49] to segment objects in each generated frame and compute their CLIP similarity with a reference image (Obj. Sim.). Due to motion and occlusion in the wrist-mounted (left and right) views, object tracking becomes unreliable. Therefore, we evaluate this metric only on the center view to ensure accuracy. All the above metrics are computed on a per-frame basis, and we report the average value across all frames as the final score. Additionally, we compute the FVD [55] between the generated and real videos to further assess the overall realism and temporal coherence of the generated data.
Geometry Consistency Analysis. We first validate the geometric consistency in RoboTransfer’s generative capability. As shown in Table 1, directly using raw depth sensor data (D.S.) as a conditional input introduces considerable noise, which adversely affects the generation quality. This degradation is evident across all three views in terms of depth accuracy, surface normals, and multi-view consistency. In contrast, employing model-predicted [9] depth (D.P.) results in smoother depth maps, which serve as a more stable supervisory signal for training the generative model. Consequently, the generated videos exhibit significantly better geometric consistency. Specifically, compared to D.S.-based generation, D.P.-based conditions relatively improves the RMSE and Mean Err. of the middle view by 27.4% and 14.2%, and increases Pix.Mat. of the middle and left/right views by 4.7% and 37.6%, respectively. Further improvements are observed where we incorporate scaled predicted depths (Scaled D.P.) to enforce a consistent depth scale across multiple views. This multi-view scale alignment further enhances geometric consistency, leading to relative improvements of 22.4% and 16.8% in Pix.Mat. of left and right views. Finally, combining normal maps with depth as joint conditions (Scaled D.P.+N.) provides complementary geometric cues, yielding the best overall performance across all geometric metrics.
Appearance Consistency Analysis.
Table 2: Experiment comparison on different appearance conditions of RoboTransfer.
Bg. Inpaint Obj. Split Camera BG. Sim. $\uparrow$ Obj. Sim. $\uparrow$ RMSE $\downarrow$ Med.Err. $\downarrow$ Pix.Mat. $\uparrow$ FVD $\downarrow$
left 0.796 0.051 1.94 191.59 117.25
left 0.802 0.049 1.92 198.32 119.65
left 0.797 0.050 1.93 196.85 118.22
left 0.805 0.047 1.92 202.03 107.43
middle 0.712 0.847 0.137 1.46 108.65
middle 0.719 0.845 0.135 1.41 105.66
middle 0.712 0.855 0.136 1.42 105.54
middle 0.720 0.858 0.133 1.39 101.17
right 0.753 0.064 2.12 71.71 226.84
right 0.762 0.060 2.04 72.44 223.91
right 0.754 0.061 2.05 72.68 222.11
right 0.764 0.059 2.00 75.67 220.12
We then evaluate RoboTransfer’s capability in controlling visual appearance. As shown in Table 2 (row-1 and row-2 of all three camera views), we observe that directly conditioning on background images without inpainting can corrupt the geometric cues. In contrast, inpainting the background before conditioning not only preserves the geometric structure but also improves appearance consistency, leading to $\sim 1\%$ relative improvements in RMSE, median error, and background similarity. Comparing row-1 and row-3 of the middle camera view, we examine object-level appearance control. In row-1, the entire image with all the objects is encoded using a global CLIP feature. In contrast, row-3 applies a finer-grained approach by masking and encoding each object individually, followed by feature concatenation. This object-wise conditions leads to a 1% improvement in object CLIP similarity, indicating better control over individual object appearances. Finally, row-4 of all three camera views demonstrate that combining background inpainting with object-wise CLIP encoding achieves the best performance in both foreground and background appearance control. This setup yields the highest CLIP similarity scores for both regions, confirming the effectiveness of jointly modeling structured and unstructured visual components.
Visualizations. We demonstrate that RoboTransfer enables controllable manipulation of both foreground and background appearance while preserving geometric and multi-view consistency across camera views. As shown in Figure 4, RoboTransfer successfully modifies the background appearance while retaining the foreground object’s texture and scene geometry. Figure 5 illustrates that the model can alter the foreground appearance without affecting the background or spatial layout. Finally, Figure 6 demonstrates the model’s ability to jointly edit both foreground and background elements. All visualizations confirm that RoboTransfer maintains geometric and multi-view consistency throughout the generated scenes.
[…]
Figure 4: Visualizations of RoboTransfer with different background reference images.
[…]
Figure 5: Visualizations of RoboTransfer with different object reference images.
[…]
Figure 6: Visualizations of RoboTransfer with different background and object reference images.
4.2 Real Robot Experiments
To evaluate RoboTransfer, we conducted experiments on a real-world robotic platform, assessing how synthetic data impacts visual policy generalization across varied environments.
Experimental Platform and Setup Details. Our experiments were conducted on the Agilex Cobot Magic platform, equipped with two PIPER robotic arms and three Intel RealSense D435i cameras—two mounted on the wrists and one positioned overhead. We used only RGB data for policy learning, creating a challenging visual understanding problem while reflecting realistic deployment constraints.
Benchmark Task and Evaluation Methodology. We selected a dual-arm manipulation sequence with four phases: (i) left arm grasps a spoon; (ii) spoon is placed on a tray; (iii) right arm grasps the spoon; (iv) spoon is placed in a cup. We evaluated across three test conditions: identical objects/background (Same), novel object appearances (Diff-Obj), and variations in both objects and environment (Diff-All). Testing protocols are detailed in the supplementary material. We introduced a Stage Score (0-4) corresponding to successfully completed phases, providing finer-grained insights than binary success/failure metrics. This reveals how far policies progress through the task sequence, particularly valuable when analyzing generalization.
Implementation Details. We adopted ACT [67] as our baseline architecture without modifications. For each task, we collected 100 expert demonstrations via ALOHA [68] teleoperation. Our methodology follows a pretrain+finetune paradigm. We trained baseline policies on real data for 100k steps, then fine-tuned with synthetic data for 50k additional steps (learning rate reduced to 1/10). For the grasp spoon task, training on real data required 8 NVIDIA H20 GPUs for approximately 24 hours, while the synthetic data fine-tuning phase utilized the same hardware configuration for about 12 hours. Additional details of the training protocol are given in the supplementary material.
Effectiveness of Synthetic Data Augmentation.
Table 3: Performance comparison across different test conditions and models. Success Rate (%) and Stage Score (0-4) are reported for each configuration. The best results for each test conditions are highlighted in bold.
Model Same Diff-Obj Diff-All
Success (%) Score Success (%) Score Success (%) Score
Baseline (Real only) 80.0 3.6 33.3 2.7 13.3 1.6
+ Obj Augmentation 44.4 3.0 22.2 2.0
+ Obj+Bg Augmentation 46.7 3.0
Table 3 presents the comparative results. Results show significant performance degradation when baseline policies face novel visual contexts, with success rates dropping from 80.0% to 13.3%. However, policies fine-tuned with our generated data show considerable resilience. In Diff-Obj conditions, success rates improved by $33.3\%$ (relative), while in Diff-All scenarios, we observed a $251\%$ relative improvement. Stage Score results further confirm that augmented models make more progress through task phases, indicating more robust manipulation understanding even when tasks aren’t fully completed.
Effect of Synthetic Data Proportions. We investigated how different proportions of synthetic data affect performance by evaluating synthetic ratios of 25%, 50%, 75%, and 100%. Figure 7 shows these results for the Diff-All conditions. Performance follows a non-linear relationship with synthetic data proportion, peaking at 50% and showing diminishing returns beyond. Success rates increased from 13.3% with no synthetic data to 46.7% at 50% synthetic ratio—a 3.5x improvement. Stage scores similarly peaked at 3.0 with 50% mixture versus 1.6 without synthetic data. Notably, using exclusively synthetic data (100%) yielded 40.0% success rate and 2.9 stage score—improvements over purely real data but below the balanced 50% mixture. This confirms complementary contributions: synthetic data provides diversity while real data contributes physical fidelity, with optimal interaction at equal proportions. RoboTransfer preserves geometric consistency while enabling systematic appearance variation, addressing a fundamental challenge in robot learning. Additional details on model configuration and extended evaluations are in the supplementary material.
Figure 7: Performance metrics across synthetic data mixing ratios. Left: Success rate vs. ratio. Right: Average stage score vs. ratio. Note that 0% represents our baseline model trained only on real data. Both metrics peak at 50% synthetic data, suggesting an optimal balance between synthetic diversity and real-world physical fidelity.
5 Conclusion
In this work, we proposed RoboTransfer, a novel diffusion-based data synthesis framework in robotics, which integrates multi-view geometry while offering explicit control over background and object attributes. We also introduced our dataset processing pipeline, which provides triplets incorporating both global geometry and appearance control conditions. The experiments and evaluations show that our RoboTransfer can produce multi-view consistency data for training robotic manipulation policies. Therefore, the generalization of visuomotor policy gets significantly improved.
Limitation. Although the method is general and flexible, several limitations remain that we aim to address in future work. First, our approach currently focuses on augmenting the appearance of objects and scenes, rather than generating new motions. Moving forward, we plan to integrate RoboTransfer with additional simulators to broaden its applicability. Furthermore, we aim to scale up both the size and diversity of the training data, which will enhance the variety and robustness of the generated outputs.
Appendix
[…]
1
Appendix A Dataset Construction Details
For the training of RoboTransfer, we collected data and designed a dedicated dataset construction pipeline (see Figure 3). In this section, we provide additional implementation details.
A.1 Data Collection
Robot Platform: We construct a large-scale robotic demonstration dataset using the Agilex Cobot Magic platform, adhering to standardized data collection protocols established in prior work [44, 39]. The robot is equipped with two PIPER robotic arms. As illustrated in Figure 8, each demonstration includes synchronized RGB-D streams captured from three Intel RealSense D435i cameras: one mounted at two hand-eye viewpoint and one positioned overhead.
Dataset Design: To capture a broad spectrum of manipulation scenarios, we systematically vary tasks, backgrounds, and manipulable objects. Specifically, we design twelve distinct manipulation tasks (visualized in Figure 9), each representing diverse interaction patterns. For each task, we collect 100 demonstration segments across 10 unique object configurations, resulting in a total of 1,000 samples per task. The dataset incorporates a variety of backgrounds, including textured tabletops and cluttered surfaces, to reflect realistic operating conditions. Manipulable objects are diversified in terms of shape, size, and material, enhancing both the visual richness and physical complexity of the dataset. This comprehensive design ensures a high degree of diversity and realism, which is crucial for training robust and generalizable visual policies.
[…]
Figure 8: Visualizations of robot platform.
[…]
Figure 9: Visualizations of the dataset.
A.2 Geometry conditions
Depth conditions: To enforce both spatial and temporal consistency, we incorporate geometry-conditioned supervision into our dataset construction pipeline. During data collection, we obtain synchronized RGB-D streams, including multi-view depth maps. However, depth maps from the RGB-D sensors are often noisy and incomplete, particularly in regions with low reflectivity (e.g., black surfaces), leading to domain gaps across samples. In contrast, depth maps rendered from simulators are typically complete and noise-free.
While we explored existing depth completion methods, we found that many suffer from limited temporal consistency and poor generalization across diverse scenes. To address these limitations, we adopt the state-of-the-art Video Depth Anything (VDA) model [9], which produces temporally coherent and spatially complete depth maps.
Despite the high-quality temporal consistency of depth estimates from Video Depth Anything (VDA), these predictions lack a globally consistent scale across multiple views. To address this, we align the VDA-predicted depth maps with the RGB-D sensor measurements via a robust least-squares scale fitting procedure. While standard least-squares regression offers a closed-form solution for affine scale alignment, it is susceptible to errors introduced by sensor noise and depth outliers. To mitigate this, we introduce a robust multi-frame fitting strategy (Figure 10) that aggregates depth correspondences across time and iteratively filters outliers based on residual error thresholds. This iterative refinement ensures stable and accurate recovery of the global scale transformation. Our approach effectively preserves the geometric completeness provided by VDA while anchoring depth predictions to the real-world scale captured by the RGB-D sensor. The resulting alignment ensures both spatial accuracy across views and temporal smoothness across frames, enabling reliable depth conditions.
Input:
Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
Output:
Scaled Depths $\mathbf{D}_{\text{scaled}}\in\mathbb{R}^{B\times H\times W}$,
1
2Initialize Mask: $\mathcal{M}\leftarrow(\mathbf{D}_{\text{sensor}}>\epsilon)\land(\mathbf{D}_{\text{pred}}>\epsilon)$ ;
3
4for 2 iterations do
5       $\mathcal{E}\leftarrow|\mathbf{D}_{\text{pred}}-\mathbf{D}_{\text{sensor}}|\odot\mathcal{M}$ ;
6       $\tau\leftarrow\mathcal{P}_{80}(\mathcal{E}[\mathcal{M}>0])$ ;
7       $\mathcal{M}\leftarrow(\mathcal{E}<\tau)\odot\mathcal{M}$ ;
8       $s,b\leftarrow\textbf{Scale Fitting}(\mathbf{D}_{\text{pred}},\mathbf{D}_{\text{sensor}},\mathcal{M})$ ;
9       $\mathbf{D}_{\text{scaled}}\leftarrow s\cdot\mathbf{D}_{\text{pred}}+b$ ;
10
return $\mathbf{D}_{\text{scaled}}$;
Algorithm 1 Dynamic Mask Refinement
Input:
Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
Mask $\mathcal{M}\in\{0,1\}^{B\times H\times W}$
Output: Scale $s\in\mathbb{R}$, Shift $b\in\mathbb{R}$
1 Flatten and Mask:
$\mathbf{p}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{pred}}\odot\mathcal{M})\in\mathbb{R}^{N}$
$\mathbf{s}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{sensor}}\odot\mathcal{M})$
[…]
2Construct Global System:
$\begin{bmatrix}\mathbf{p}^{\top}\mathbf{p}&\mathbf{p}^{\top}\mathbf{1}\\ \mathbf{p}^{\top}\mathbf{1}&\mathbf{1}^{\top}\mathbf{1}\end{bmatrix}\begin{bmatrix}s\\ b\end{bmatrix}=\begin{bmatrix}\mathbf{p}^{\top}\mathbf{s}\\ \mathbf{1}^{\top}\mathbf{s}\end{bmatrix}$
Closed-Form Solution:
$s=\frac{(\mathbf{1}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{1}^{\top}\mathbf{s})}{\det}$
$b=\frac{(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})}{\det}$
where $\det=(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{1})-(\mathbf{p}^{\top}\mathbf{1})^{2}>\epsilon$
Algorithm 2 Scale Fitting
Figure 10: Depth Scale Alignment: Implementation Details
Normal conditions: For high-precision surface reconstruction, surface normals are critical and can provide surface details rather than only depth. At the same time, normal does not have scale consistency issues. We compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS [24].
A.3 Appearance conditions
Object Mask: For the manipulation task data, we first select keyframes that effectively capture both the objects and the background conditions of the scene. We then extract all operable objects using a structured prompt template (see Figure 11), which describes each object in terms of its color, material, shape, and spatial position. These object descriptions are fed into Grounding DINO[38] to obtain bounding boxes for object detection. The resulting bounding boxes are subsequently used as prompts for SAM2[49] to generate the corresponding object masks.
*System Instruction*
As a visual descriptor, articulate objects with this flow:
Color -> Material -> Shape/Texture -> Spatial Position
*Template Format*
"A [color] [material] [object], [shape] with [texture], located [region](x,y)"
*Demonstration Examples*
- A red plastic ball, spherical with a smooth glossy surface, at the center (250,300)
- A brown wooden chair, angular with a matte wood grain, in top-left (100,50)
- A silver-blue metallic arm, mechanical with a brushed texture, on the right (600,200)
Figure 11: Visual Description Prompt Template Architecture.
Objects and Background conditions:To extract the object conditions, we first segment image patches corresponding to individual objects from the full image. Each patch is resized to 384×384 and passed through the CLIP model [47] to obtain RGB embeddings. For the background conditions, we mask out all detected objects from the full image using their corresponding masks. The masked image is then fed into an inpainting model to reconstruct the background with the objects removed.
Appendix B Visual Policy Implementation Details
To validate the effectiveness of the data generated by RoboTransfer, we train a visual policy using the procedures detailed below for both training and deployment.
B.1 Data Collection and Preprocessing
To ensure the validity of our evaluation, we excluded the dataset described in Section A from the real-robot experiments. The data preparation proceeded as follows:
Real Expert Data: We collected 100 sets of expert demonstration data for each manipulation task using the ALOHA teleoperation system. The primary observations were RGB images (1280x720 resolution, subsequently downscaled to 640$\times$360 to optimize training), with camera streams captured at approximately 30Hz and sampled at 10Hz for policy training. Auxiliary robot state information (e.g., joint positions, end-effector poses) was originally recorded at 200Hz and downsampled to 50Hz for model use.
Synthetic Data Generation: To further enhance model performance and generalization, we generated synthetic data based on the collected real data, specifically targeting variations in foreground objects and background scenes, for use in the subsequent fine-tuning phase.
B.2 Training Pipeline and Core Parameters
Our training pipeline is based on the ACT (Action Chunking Transformer) architecture, which processes visual input from three distinct perspectives (two wrist-mounted cameras and one fixed overhead camera). At each training timestep, the model receives a single RGB frame from each view.
Training Objective and Two-Stage Training Strategy: The policy model is designed to predict a sequence of the next 100 robot states, corresponding to a 2-second action horizon given the 50Hz robot state frequency.
We adopted a "pre-train-then-finetune" paradigm:
Pretraining on Real Data:The model was first pretrained for 100k steps using the collected real expert demonstration data. During this phase, the batch size was set to 512, and the learning rate was $1\times 10^{-4}$.
Finetuning with Synthetic Data: After pretraining, the aforementioned synthetic data was introduced to fine-tune the model for an additional 50k steps. The learning rate for this phase was reduced to $1\times 10^{-5}$.
All training was performed on a cluster equipped with 8 NVIDIA H20 GPUs. The pretraining phase took approximately 24 hours, and the fine-tuning phase required about 12 hours.
B.3 Real-Robot Deployment and Evaluation
The trained policies were deployed and evaluated in the following environment:
Deployment Platform: Experiments were conducted on the Agilex Cobot Magic platform8, the same platform used for initial data collection, ensuring consistency between training and testing environments.
Inference Procedure: During deployment, the policy operated in a synchronous execution mode. Specifically, the system waits for the robot to fully execute the action sequence (comprising 100 actions) generated from the model’s previous inference step before capturing new observations and proceeding to the next model inference and subsequent execution. At each decision point, the model receives as input a single RGB frame from each of its multiple camera views and the current robot state and outputs the sequence of 100 actions. The latency for this single model inference process is approximately 10ms. These actions are then executed on the robot at a 50Hz control frequency, consistent with the robot state prediction rate during training.
Appendix C More Qualitative Results
C.1 Real-to-Real Transfer
The proposed RoboTransfer offers strong controllability over both foreground and background generation. Given the same structured input, the model allows for flexible editing of background attributes such as texture and color, as shown in Figure 12. Meanwhile, as illustrated in Figure 13, the appearance of foreground objects, including their color, can also be effectively modified. This real-to-real generation framework enriches the diversity of training data for downstream tasks, contributing to better generalization of the policy model.
[…]
Figure 12: Visualizations of RoboTransfer with different background reference images.
[…]
Figure 13: Visualizations of RoboTransfer with different object reference images.
C.2 Sim-to-Real Transfer
As shown in Figure 14, RoboTransfer generates photorealistic videos from simulated structural inputs, including out-of-distribution cases. This Sim-to-Real paradigm reduces dependency on structured information from real-world datasets, thereby better supporting downstream robotic learning tasks.
[…]
Figure 14: Visualizations of Sim-to-Real generation results.
C.3 Comparison
Figure 15 illustrates that Cosmos Transfer [2] is mainly effective for robot scene generation under fixed camera settings (central view), while its quality drops notably for dynamic camera viewpoints (side views). Compared to Cosmos Transfer [2], RoboTransfer demonstrates significantly better multi-view consistency, enabling more realistic and coherent novel-view synthesis across varying camera poses.
[…]
Figure 15: Comparison with Cosmos generation results.

### v2

@@
-RoboTransfer: Geometry-Consistent Video Diffusion for Robotic Visual Policy Transfer
+RoboTransfer: Controllable Geometry-Consistent Video Diffusion for Manipulation Policy Transfer
 Abstract
-Imitation Learning has become a fundamental approach in robotic manipulation.
-However, collecting large-scale real-world robot demonstrations is prohibitively expensive.
-Simulators offer a cost-effective alternative, but the sim-to-real gap make it extremely challenging to scale.
-Therefore, we introduce RoboTransfer, a diffusion-based video generation framework for robotic data synthesis.
-Unlike previous methods, RoboTransfer integrates multi-view geometry with explicit control over scene components, such as background and object attributes.
-By incorporating cross-view feature interactions and global depth/normal conditions, RoboTransfer ensures geometry consistency across views.
-This framework allows fine-grained control, including background edits and object swaps.
-Experiments demonstrate that RoboTransfer is capable of generating multi-view videos with enhanced geometric consistency and visual fidelity.
-In addition, policies trained on the data generated by RoboTransfer achieve a 33.3% relative improvement in the success rate in the Diff-Obj setting and a substantial 251% relative improvement in the more challenging Diff-All scenario.
-Explore more demos on our project page: https://horizonrobotics.github.io/robot_lab/robotransfer
-**footnotetext: Equal contribution. footnotetext: Corresponding author.
+The goal of general-purpose robotics is to create agents that can seamlessly adapt to and operate in diverse, unstructured human environments.
+Imitation learning has become a key paradigm for robotic manipulation, yet collecting large-scale and diverse demonstrations is prohibitively expensive.
+Simulators provide a cost-effective alternative, but the sim-to-real gap remains a major obstacle to scalability.
+We present RoboTransfer, a diffusion-based video generation framework for synthesizing robotic data.
+By leveraging cross-view feature interactions and globally consistent 3D geometry, RoboTransfer ensures multi-view geometric consistency while enabling fine-grained control over scene elements, such as background editing and object replacement.
+Extensive experiments demonstrate that RoboTransfer produces videos with superior geometric consistency and visual fidelity.
+Furthermore, policies trained on this synthetic data exhibit enhanced generalization to novel, unseen scenarios.
+Project page: https://horizonrobotics.github.io/robot_lab/robotransfer.
+[…]
+Figure 1: Overview of RoboTransfer.
+While real-world data collection is prohibitively expensive and simulated data often lacks visual fidelity, RoboTransfer synthesizes high-quality data with strict multi-view consistency.
+Our experiments demonstrate that this synthetic data significantly enhances policy performance on real robots.
+11footnotetext: Equal contribution.22footnotetext: Corresponding author: nemo.liu@horizon.auto, zhizhong.su@horizon.auto
+Table 1: Comparison of different methods for robotic data generation. RoboTransfer is the only one that provides temporal and multi-view consistency while offering fine-grained control.
+Method Model Type Generation Consistency Control Level
+Temporal Multi-View Background Object Environment
+ROSIE [54] Image Diffusion
+RoboEngine [55] Image Diffusion
+Cosmos-Transfer1 [2] Video Diffusion
+RoboTransfer (Ours) Video Diffusion
 1 Introduction
-Imitation Learning (IL) has become a fundamental approach for visuomotor control in robotic manipulation [67, 68].
+Imitation Learning (IL) has become a fundamental approach for visuomotor control in robotic manipulation [60].
 However, collecting large-scale real-world robot demonstrations is prohibitively expensive [6, 7].
-Simulated environments offer a cost-effective alternative [54, 59, 19, 44], but the scarcity of assets and sim-to-real gap[51, 43] make it extremely challenging to scale.
+Simulated environments offer a cost-effective alternative [51, 37], but the scarcity of assets and sim-to-real gap[43, 36] make it extremely challenging to scale.
 Recently, diffusion models [21] have gained attention as a promising method to synthetically generate realistic and diverse data.
@@
 In contrast, Cosmos-Transfer [2] generates photorealistic data using a video diffusion model conditioned on segmentation and depth, which helps preserve geometric consistency.
-Despite the rapid advancements in the field, two key challenges remain.
-First, robotic systems often rely on observations from a single viewpoint, which inherently limits their perception.
-To overcome this limitation, multi-view observation is a commonly used approach.
-However, generating multi-view consistent results remains a significant challenge for video generative models.
-Second, robot manipulation scenarios are typically complex and highly interactive.
-Controlling such scenarios precisely through textual input alone, as demonstrated in text-to-video frameworks [5, 70], presents substantial difficulties.
-In this work, we introduce RoboTransfer, as is shown in Figure 1, a video generation framework designed for data synthesis in robotics.
-To the best of our knowledge, it is the first framework to integrate multi-view geometry while offering explicit control over background, object attributes.
-RoboTransfer incorporates cross-view feature interactions and global depth/normal conditions within the video diffusion architecture, ensuring 3D consistency across different camera views and over time.
-To enhance controllability, RoboTransfer decomposes each scene into distinct control conditions, enabling background edits and object swaps by conditioning on this decomposition.
-[…]
-Figure 1: Overview of RoboTransfer.
-Our work makes the following contributions:
+Despite rapid advancements, two key challenges persist.
+First, robotic systems often rely on single-view observations, limiting their perception.
+Multi-view observation is commonly used to address this, but generating consistent multi-view results remains a significant challenge for video generative models.
+Second, robot manipulation tasks are complex and interactive, and precisely controlling these tasks via textual input alone, as seen in text-to-video frameworks [5, 63], remains a substantial challenge.
+To address these challenges, we introduce RoboTransfer, a geometry-consistent video diffusion framework tailored for robotic visual policy transfer (Figure 1).
+As summarized in Table 1, RoboTransfer is, to the best of our knowledge, the first framework to guarantee multi-view generation consistency in robotic data synthesis, while also offering fine-grained and disentangled control over both background and object attributes.
+The main contributions are as follows:
   1. 1.
-    We propose RoboTransfer, a robot generation framework based on multi-view, geometry-aware video diffusion models that maintain fidelity across multiple reference conditions.
+    We propose RoboTransfer, a video data generation framework for robotic manipulation that ensures multi-view consistency while enabling fine-grained, disentangled control.
   2. 2.
-    We introduce a dataset construction pipeline that decomposes real-world recording data into high-quality triplets, which incorporate global geometry and appearance control conditions.
+    We introduce a novel data construction pipeline that automatically decomposes real-world robot demonstrations into the geometric and appearance conditions.
   3. 3.
-    Through carefully designed experiments and evaluations, we demonstrate that our approach generates multi-view videos with significantly improved geometric consistency and visual fidelity.
-Furthermore, policies trained with RoboTransfer-generated data exhibit better generalization to novel and unseen environments.
+    We demonstrate that RoboTransfer generates multi-view videos with substantially improved generation consistency, and that policies trained on this data generalize more effectively to novel, unseen environments.
 2 Related Work
 2.1 Visual Generalizable Imitation Learning
-Imitation Learning (IL) has emerged as a foundational approach for visuomotor control in robotic manipulation.
-Deep networks trained on collected demonstrations—e.g., DP [12], OpenVLA [31], Act [67], Aloha [68], E2E-VP [36], and RT2 [7]—map raw visual inputs directly to motor commands.
-To bolster policy robustness and generalization, researchers have explored novel architectures [12] and vision pretraining schemes [4, 13].
-However, gathering large-scale, task-specific robot data remains prohibitively expensive.
-Consequently, many methods leverage auxiliary sources—human demonstration videos [20], high-fidelity simulators [54, 59, 19, 44], or operational logs from other platforms [46, 30].
-Each of these domains exhibits its own biases and visual characteristics, introducing distributional shifts that can degrade IL policy performance when transferred to the target environment.
-In this work, we offer a complementary perspective on generalizable visual imitation learning by leveraging state-of-the-art video diffusion models.
-By generating target-conditioned training data, our approach significantly enhances the robustness of policy and enables seamless transfer to novel environments.
-2.2 Data Augmentation for Policy Learning
-Data augmentation is widely used to train robust machine learning models.
-Prior work has applied augmentation directly to real-world robot captures to improve the resilience of visual policies against environmental variations[41, 42, 15, 23, 21].
-Image-inpainting methods further enhance appearance invariance by replacing or masking out scene elements in collected frames [62, 11, 10, 40].
-Likewise, neural view synthesis techniques generate novel viewpoints from existing robot data, promoting view-invariant feature learning [53, 8].
-Simulators can cheaply produce vast quantities of labeled frames [54, 59, 19, 44], but discrepancies in physics modeling, rendering fidelity, and scene composition often hinder direct sim-to-real transfer.
-Domain randomization is used to close the sim-to-real gap, thereby enabling policies trained in simulation to generalize more effectively to real-world environments.
-[35, 22, 34, 51, 1, 48, 25].
-In contrast to these methods, we leverage a video diffusion model to generate photorealistic image manipulations.
-This approach preserves both multi-view and geometry consistency, producing richly detailed samples for robot policy training.
+Imitation Learning (IL) has become a cornerstone for visuomotor control in robotic manipulation, with deep networks mapping raw visual inputs to motor commands based on demonstrations [10, 26, 60, 7].
+However, collecting large-scale task-specific data is prohibitively expensive, leading to the use of auxiliary sources such as human demonstration videos [17, 4] and operational logs from other platforms [39, 25, 8].
+These sources introduce domain-specific biases and distributional shifts, which can degrade IL policy performance when transferred to new environments.
+While simulators can generate large quantities of labeled frames [45, 51, 16, 37], discrepancies in physics modeling, rendering fidelity, and scene composition hinder sim-to-real transfer.
+Domain randomization attempts to close this gap [30, 18, 29, 1], but it typically only applies color perturbations across the entire image, without capturing localized or structurally meaningful variations.
+In this work, we leverage a generative model to synthesize photorealistic multi-view videos for manipulations, significantly enhancing policy robustness and enabling seamless transfer to novel environments.
+[…]
+Figure 2: The RoboTransfer framework performs multi-view consistent modeling to jointly reason across viewpoints.
+It represents geometry with metric depth and normal maps, and encodes appearance using reference backgrounds and object images for detailed control over appearance.
 2.2 Generation Models in Embodied AI
-With the rapid advancement of generative models such as HunyuanVideo [33], Wan [56], SVD [5], Cosmos [2], CogVideoX [61], OpenSora [70], there is a growing trend toward leveraging generated data to bridge the gap between synthetic environments and real-world physical applications [72].
+With the rapid advancement of generative models  [28, 47, 5, 2, 53, 63, 13, 56, 31, 19], there is a growing trend toward leveraging generated data to bridge the gap between synthetic environments and real-world physical applications [65].
 In particular, domains like autonomous driving have seen increasing adoption of generative techniques, as demonstrated by [49, 59, 22, 50, 14, 15, 57, 38, 58].
-Similarly, robotic scenarios—where real-world data collection is often prohibitively expensive and time-consuming—have greatly benefited from generative data.
+Similarly, in robotic scenarios, where real-world data collection is often prohibitively expensive and time-consuming, generative data has proven to be highly beneficial.
 Models such as UniSim [52], UniPi [11], and RoboDreamer [64] synthesize future robot behaviors via text prompts, which are then translated into actionable commands using inverse dynamics models.
@@
 Finally, the dataset construction pipeline is elaborated in Sec.
-3.2.
+3.3.
 3.1 Video Diffusion Model Preliminaries
-Diffusion-based methods for controllable video generation typically model the process as a gradual refinement of a noise latent $\boldsymbol{\epsilon}$ into a clean video latent $\mathbf{x_{0}}$, under the guidance of spatially aligned conditions $y_{s}$ (e.g., depth map, segmentation map) and unstructured conditions $y_{u}$ (e.g., CLIP embedding).
+Diffusion-based methods for controllable video generation model the process as a gradual refinement of a noise latent $\boldsymbol{\epsilon}$ into a clean video latent $\mathbf{x_{0}}$, under the guidance of spatially aligned conditions $y_{s}$ (e.g., depth map) and unstructured conditions $y_{u}$ (e.g., CLIP embedding).
 Popular approaches in this paradigm include Flow Matching [32], DDPM [21], and EDM [24].
 In the case of EDM, the learning objective is formulated as:
-$\mathcal{L}(D_{\theta},\sigma)=\mathbb{E}_{\mathbf{x_{0}},y_{s},y_{u}}\left[\left\|\mathbf{x_{0}}-D_{\theta}\left(\mathcal{E}(\mathbf{x_{0}}+\mathbf{n}),\tau_{u}(y_{u}),\tau_{s}(y_{s}),\sigma\right)\right\|_{2}^{2}\right].$ (1)
+$\mathcal{L}(D_{\theta},\sigma)=\mathbb{E}_{\mathbf{x_{0}},y_{s},y_{u}}\left[\left\|\mathbf{x_{0}}-D_{\theta}\left(\mathcal{E}(\mathbf{x_{0}}+\mathbf{n}),\tau_{u}(y_{u}),\tau_{s}(y_{s}),\sigma\right)\right\|_{2}^{2}\right].$
 Here, $\mathbf{x_{0}}\sim p_{\text{data}}$ denotes a clean sample drawn from the dataset, and $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ represents Gaussian noise.
@@
 To encourage stable learning across varying noise magnitudes, the total training objective aggregates the per-noise loss via a weighted expectation:
-$\mathcal{L}(D_{\theta})$ $=\mathbb{E}_{\sigma}\left[\frac{\lambda(\sigma)}{\exp(u(\sigma))}\mathcal{L}(D_{\theta},\sigma)+u(\sigma)\right],$ (2)
+$\vskip-2.0pt\mathcal{L}(D_{\theta})$ $=\mathbb{E}_{\sigma}\left[\frac{\lambda(\sigma)}{\exp(u(\sigma))}\mathcal{L}(D_{\theta},\sigma)+u(\sigma)\right],$ (2)
 $\lambda(\sigma)$ $=\frac{\sigma^{2}+\sigma_{\text{data}}^{2}}{(\sigma\cdot\sigma_{\text{data}})^{2}},$ (3)
-$\ln(\sigma)$ $\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2}).$ (4)
+$\ln(\sigma)$ $\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2}).\vskip-2.0pt$ (4)
 In this setup, $\sigma_{\text{data}}$ denotes the empirical standard deviation of the data, while the distribution of $\sigma$ is governed by the hyperparameters $P_{\text{mean}}$ and $P_{\text{std}}$.
 The weighting function $\lambda(\sigma)$ ensures that all noise levels contribute proportionally during the initial stages of training.
+3.2 RoboTransfer Framework
 […]
-Figure 2: The RoboTransfer framework performs multi-view consistent modeling to jointly reason across viewpoints.
-It represents geometry with scaled depth and normal maps, and encodes appearance using reference backgrounds and object-specific images for detailed control over appearance.
-3.2 RoboTransfer Framework
+Figure 3: RoboTransfer data construction pipeline generates image pairs with Geometry and Appearance conditions.
+Geometry conditions (left) with metric depth and normals from demonstration videos and appearance conditions (right) from keyframes.
+A VLM-based descriptor generates object descriptions, which are processed by Grounding-SAM to create per-object masks.
 The overall framework of RoboTransfer is illustrated in Figure 2.
@@
 On the conditions side, RoboTransfer incorporates fine-grained control by encoding both geometric and appearance information.
-Specifically, we represent geometry using scaled depth maps and surface normal maps, capturing the underlying 3D structure of the scene.
+Specifically, we represent geometry using metric depth maps and surface normal maps, capturing the underlying 3D structure of the scene.
 Meanwhile, the appearance is encoded using reference background images and object-specific images, providing detailed control over texture, color, and contextual appearance.
-In the following sections, we first introduce the multi-view consistent modeling, and then describe the encoding mechanisms for geometric and appearance conditions, respectively.
+In the following sections, we first introduce the multi-view consistent modeling and then describe the encoding mechanisms for geometric and appearance conditions, respectively.
 Multi-view Consistent Modeling.
-In robotic environments, it is common to deploy multiple cameras from different viewpoints to capture a scene simultaneously.
+Robotic manipulation often relies on multi-view camera setups to capture a scene in parallel.
 To generate multi-view consistent videos for downstream tasks, RoboTransfer leverages the multi-view in-context learning capabilities [59, 23] inherently present in pretrained diffusion models.
-Specifically, given $N$ synchronized videos $\{V_{1},V_{2},...,V_{N}\}$ from different viewpoints, we concatenate them along the width dimension and encode the joint information using a VAE encoder $\mathcal{E}$:
-$\mathbf{x}_{0}=\mathcal{E}([V_{1},V_{2},\ldots,V_{N}]),$ (5)
-where $[V_{1},V_{2},\ldots,V_{N}]$ represents the concatenated multi-view video sequence.
-This modeling approach maintains the integrity of the pretrained video diffusion architecture, requiring no structural modifications, while enabling fast convergence and high-quality, view-consistent video generation in robotic settings.
-Geometry Conditions.
+Specifically, given $N$ synchronized videos $\{V_{1},V_{2},...,V_{N}\}$ from different viewpoints, we concatenate them along the width dimension and encode the joint information using a Variational Autoencoder(VAE) encoder $\mathcal{E}$:
+$\mathbf{x}_{0}=\mathcal{E}([V_{1},V_{2},\ldots,V_{N}]),\vskip-2.0pt$ (5)
+where $V_{1},V_{2},\ldots,V_{N}]$ represents the concatenated multi-view video sequence.
+This modeling approach leverages the spatial reasoning capabilities of existing video diffusion backbones, requiring no structural modifications, while enabling fast convergence and high-quality, view-consistent video generation in robotic manipulation settings.
+Geometry Conditions Injection.
 Recent video generation models primarily learn spatiotemporal coherence through data-driven approaches such as temporal attention [5, 63] or full attention [47, 28, 2].
@@
 Specifically, we utilize depth and surface normal videos as geometric conditions.
-These two types of geometric cues are complementary — depth maps provide information about spatial distances from the camera, while normal maps encode local surface orientations.
+These two types of geometric cues are complementary; depth maps provide information about spatial distances from the camera, while normal maps encode local surface orientations.
 When combined, they offer a richer and more holistic description of scene geometry (see Sec.
 3.3 for details on data acquisition).
-Since both the depth and normal videos are spatially aligned with the RGB views, we propose GuideNet, which utilizes a stack of convolutional layers that jointly downsample and encode the depth and normal sequences.
-The resulting geometry-aware representation is then concatenated with the noise latents along the channel dimension, allowing the generation process to be guided by accurate 3D structural cues.
-Appearance Conditions.
+Since the depth and surface normal sequences are spatially aligned with their corresponding RGB views, we leverage the VAE encoder composed of stacked convolutional layers that jointly downsample and encode geometric cues from depth and normal videos.
+The resulting geometry-aware representation is concatenated with the noise latents along the channel dimension, enabling diffusion-based video generation to be precisely guided by consistent and physically plausible 3D cues throughout the generation process.
+Appearance Conditions Injection.
 To enable fine-grained control over the generated textures, RoboTransfer introduces texture conditions from two complementary perspectives: background appearance and object appearance.
-Specifically, we use a background reference image and a set of object reference images as conditions signals to guide the generation process toward faithful reproduction of scene textures and object-level details.
+Specifically, we use a background reference image and a set of object reference images as condition signals to guide the generation process toward faithful reproduction of scene textures and object-level details.
 A key challenge in appearance control is ensuring compatibility with the geometry conditions.
-Naïvely combining appearance and geometry inputs may introduce conflicts (e.g., mismatched depth and texture), thereby weakening both the structural consistency and the visual fidelity of generated results.
+Naïvely combining appearance and geometry inputs may introduce conflicts (e.g., mismatched depth and texture), thereby weakening both the geometry consistency and the visual fidelity of generated results.
 To address this, we carefully curate the appearance reference images, resulting in a background reference image $C_{b}$ and object reference images $C_{o}$ that do not contradict the geometric priors (see Sec.
@@
 This design allows flexible and scalable conditioning on diverse object appearances while preserving compatibility with the structured geometry inputs.
-[…]
-Figure 3: RoboTransfer data processing pipeline.
-The data processing pipeline consists of two main components: Geometry conditions (left) are derived by using Mono-Normal and Video Depth Anything, which are scale-aligned with sensor depth to ensure consistency.
-Appearance conditions (right) are obtained by sampling keyframes from the video.
-GPT-4 is used to generate object descriptions, which are then processed by Grounding-SAM to create per-object masks.
-Additionally, background inpainting is used to generate complete reference backgrounds.
-3.3 Dataset Construction from Real World Data
-Data construction is central to our framework.
-The overall pipeline is illustrated in Figure 3.
-Unlike conventional text-to-video or image-to-video tasks, our approach requires both globally consistent geometry and multiple part-specific reference images (e.g., background and manipulable objects).
-Our pipeline begins with the collection of a large-scale robotic demonstration dataset.
-Each demonstration captures synchronized RGB-D streams from both a hand-eye and an overhead RGB-D camera.
-To cover a wide range of operating conditions, we vary backgrounds (e.g., textured tabletops, cluttered surfaces) and manipulable objects (shapes, sizes, materials).
-To enforce spatial and temporal consistency, we incorporate geometric cues such as depth and surface normals.
-Since raw depth maps from RGB-D sensors are often noisy, we employ the state-of-the-art video depth estimator, Video Depth Anything[9], to produce temporally consistent depth maps.
-We then align the scale of the estimated depth map with the RGB-D sensor’s depth using robust least-squares fitting to ensure global spatial accuracy.
-Additional details are provided in the supplementary material.
-For surface normals, we compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS[24], which jointly serves as reliable geometric conditions.
-Next, we sample a subset of RGB keyframes as visual references.
-These frames are processed by GPT-4 to generate structured scene descriptions, including object-level descriptions for individual entities.
-These descriptions guide the Grounding-SAM module[50], which detects and segments per-object masks.
-To obtain clean background references, we inpaint the segmented objects from the original images, generating plausible object-free scenes (e.g., an empty tabletop).
-For multimodal object-level conditions, the segmented object masks are further processed with the CLIP model[47] to obtain semantic embeddings.
-These conditions allow the diffusion model to generate photorealistic, semantically grounded video augmentations that respect both object identity and spatial layout.
-4 Experiments
-4.1 Synthesis Quality
-Implementation Details.
-To construct our training dataset, we leverage the open-source Cobot Magic platform to collect a large-scale video corpus of dual-arm robot executions, covering diverse tasks such as grasping and placing.
-The raw videos are segmented into 10Hz clips of 30 frames each, resulting in approximately 24k clips for training.
-Additionally, we curate a set of 1.6k 10Hz 30-frame clips from the collected dataset for video synthesis quality evaluation.
-All videos are annotated with conditions as described in Sec.
-3.2, to facilitate both training and evaluation.
-RoboTransfer is fine-tuned from the pretrained video diffusion model [5].
-During training, videos from each camera view are resized to a resolution of $640\times 384$.
-We adopt the AdamW optimizer with a learning rate of $3\times 10^{-5}$ and a global batch size of 8, training for a total of 70K steps.
-During inference, we use the EDM scheduler [29] to perform 30 denoising steps and apply classifier-free guidance.
-Table 1: Experiment comparison on different geometry conditions of RoboTransfer.
+Table 2: Experiment comparison on different geometry conditions of RoboTransfer: It shows that combining metric depth and normal maps yields the best consistency across all views and metrics.
 Model Camera RMSE $\downarrow$ Abs.Rel. $\downarrow$ Sq.Rel. $\downarrow$ Mean Err. $\downarrow$ Med.Err. $\downarrow$ Pix.Mat. $\uparrow$ FVD $\downarrow$
@@
 RoboTransfer [D.P.] left 0.054 0.090 0.010 3.91 2.28 149.68 123.31
-RoboTransfer [Scaled D.P.] left 0.049 0.081 0.008 3.48 1.99 183.26 112.43
-RoboTransfer [Scaled D.P. + N.] left 0.047 0.079 0.008 3.31 1.92 202.03 107.43
-RoboTransfer [D.S.] middle 0.182 0.074 0.031 3.51 1.58 153.76
-RoboTransfer [D.P.] middle 0.132 0.053 0.015 3.05 1.43 95.89
-RoboTransfer [Scaled D.P.] middle 0.134 0.054 0.015 2.93 1.40 103.32
-RoboTransfer [Scaled D.P. + N.] middle 0.133 0.054 0.015 2.86 1.39 101.17
+RoboTransfer [Metric D.P.] left 0.049 0.081 0.008 3.48 1.99 183.26 112.43
+RoboTransfer [Metric D.P. + N.] left 0.047 0.079 0.008 3.31 1.92 202.03 107.43
+RoboTransfer [D.S.] head 0.182 0.074 0.031 3.51 1.58 153.76
+RoboTransfer [D.P.] head 0.132 0.053 0.015 3.05 1.43 95.89
+RoboTransfer [Metric D.P.] head 0.134 0.054 0.015 2.93 1.40 103.32
+RoboTransfer [Metric D.P. + N.] head 0.133 0.054 0.015 2.86 1.39 101.17
 RoboTransfer [D.S.] right 0.090 0.137 0.025 4.93 2.91 40.70 396.33
 RoboTransfer [D.P.] right 0.072 0.103 0.016 4.14 2.36 56.02 262.96
-RoboTransfer [Scaled D.P.] right 0.064 0.090 0.011 3.74 2.11 65.45 226.76
-RoboTransfer [Scaled D.P. + N.] right 0.058 0.087 0.009 3.55 2.00 75.67 220.12
+RoboTransfer [Metric D.P.] right 0.064 0.090 0.011 3.74 2.11 65.45 226.76
+RoboTransfer [Metric D.P. + N.] right 0.058 0.087 0.009 3.55 2.00 75.67 220.12
+Table 3: Experiment comparison on different appearance conditions of RoboTransfer.
+Bg. Inpaint Obj. Split Camera BG. Sim. $\uparrow$ Obj. Sim. $\uparrow$ RMSE $\downarrow$ Med.Err. $\downarrow$ Pix.Mat. $\uparrow$ FVD $\downarrow$
+left 0.796 0.051 1.94 191.59 117.25
+left 0.802 0.049 1.92 198.32 119.65
+left 0.797 0.050 1.93 196.85 118.22
+left 0.805 0.047 1.92 202.03 107.43
+head 0.712 0.847 0.137 1.46 108.65
+head 0.719 0.845 0.135 1.41 105.66
+head 0.712 0.855 0.136 1.42 105.54
+head 0.720 0.858 0.133 1.39 101.17
+right 0.753 0.064 2.12 71.71 226.84
+right 0.762 0.060 2.04 72.44 223.91
+right 0.754 0.061 2.05 72.68 222.11
+right 0.764 0.059 2.00 75.67 220.12
+3.3 Dataset Construction from Real Robotic Demonstration Data
+Data construction is central to our framework, where real-world recording data is decomposed into high-quality triplets for training Robotransfer, consisting of Geometry Conditions, Appearance Conditions, and ground-truth images.
+The overall pipeline is illustrated in Figure 3.
+Geometry Conditions Construction Starting from robotic demonstration data with two wrist cameras and one head camera.
+To enforce spatial and temporal consistency, we incorporate geometric cues such as depth and surface normals.
+Since some robots only have RGB sensors and raw depth maps from RGB-D sensors are often noisy, we use a state-of-the-art depth estimator [9, 48] to produce consistent depth maps.
+For unmetric depth outputs, we align the estimated scale with the RGB-D sensor using robust least-squares fitting to ensure global spatial accuracy.
+For surface normals, we compute per-frame estimates using state-of-the-art monocular normal estimators [20, 48], offering conditions for detailed geometry.
+Appearance Conditions Construction.
+To enable controllable data generation with the diffusion model, we sample multiview RGB keyframes as appearance conditions.
+A VLM-based descriptor generates scene and object descriptions, which guide the Grounding-SAM module [42] to detect and segment per-object masks.
+To obtain clean background references, the segmented objects are inpainted in the original images, producing plausible object-free scenes (e.g., an empty tabletop).
+For object-level conditions, the segmented masks are further processed with the CLIP model [40] to extract semantic embeddings.
+4 Experiments for Synthesis Quality
+[…]
+Figure 4: Visualizations of RoboTransfer with different background reference images.
+[…]
+Figure 5: Visualizations of RoboTransfer with different object reference images.
+[…]
+Figure 6: Visualizations of RoboTransfer with different background and object reference images.
+[…]
+Figure 7: Visualizations of RoboTransfer Sim-to-Real generation results.
+To evaluate the synthesis quality of RoboTransfer, including multi-view geometric consistency and controllability, we perform both quantitative and qualitative analyses.
+In this section, we first present the evaluation metrics and analysis of the synthesis quantity in Sec.
+4.1.
+Then, the synthesis quality results are presented in Sec.
+4.
+4.1 Synthesis Quantitative Analysis
+Implementation Details.
+RoboTransfer is fine-tuned from the pre-trained SVD  [5] model.
+During inference, we use the EDM scheduler [24] to perform 30 denoising steps and apply classifier-free guidance.
 Evaluation Metrics.
@@
 Consequently, the generated videos exhibit significantly better geometric consistency.
-Specifically, compared to D.S.-based generation, D.P.-based conditions relatively improves the RMSE and Mean Err.
-of the middle view by 27.4% and 14.2%, and increases Pix.Mat.
+Specifically, compared to D.S.-based generation, D.P.-based conditions relatively improve the RMSE and Mean Err.
+of the middle view by 27.4% and 14.2%, and increase Pix.Mat.
 of the middle and left/right views by 4.7% and 37.6%, respectively.
-Further improvements are observed where we incorporate scaled predicted depths (Scaled D.P.) to enforce a consistent depth scale across multiple views.
+Further improvements are observed when we incorporate metric predicted depths (Metric D.P.) to enforce a consistent depth scale across multiple views.
 This multi-view scale alignment further enhances geometric consistency, leading to relative improvements of 22.4% and 16.8% in Pix.Mat.
 of left and right views.
-Finally, combining normal maps with depth as joint conditions (Scaled D.P.+N.) provides complementary geometric cues, yielding the best overall performance across all geometric metrics.
+The Pix.Mat.
+is omitted for the head camera since it remains static.
+Finally, combining normal and depth as joint conditions (Metric D.P.+N.) provides complementary geometric cues, yielding the best overall performance across all geometric metrics.
 Appearance Consistency Analysis.
-Table 2: Experiment comparison on different appearance conditions of RoboTransfer.
-Bg. Inpaint Obj. Split Camera BG. Sim. $\uparrow$ Obj. Sim. $\uparrow$ RMSE $\downarrow$ Med.Err. $\downarrow$ Pix.Mat. $\uparrow$ FVD $\downarrow$
-left 0.796 0.051 1.94 191.59 117.25
-left 0.802 0.049 1.92 198.32 119.65
-left 0.797 0.050 1.93 196.85 118.22
-left 0.805 0.047 1.92 202.03 107.43
-middle 0.712 0.847 0.137 1.46 108.65
-middle 0.719 0.845 0.135 1.41 105.66
-middle 0.712 0.855 0.136 1.42 105.54
-middle 0.720 0.858 0.133 1.39 101.17
-right 0.753 0.064 2.12 71.71 226.84
-right 0.762 0.060 2.04 72.44 223.91
-right 0.754 0.061 2.05 72.68 222.11
-right 0.764 0.059 2.00 75.67 220.12
 We then evaluate RoboTransfer’s capability in controlling visual appearance.
@@
 In contrast, inpainting the background before conditioning not only preserves the geometric structure but also improves appearance consistency, leading to $\sim 1\%$ relative improvements in RMSE, median error, and background similarity.
-Comparing row-1 and row-3 of the middle camera view, we examine object-level appearance control.
+Comparing row-1 and row-3 of the head camera view, we examine object-level appearance control.
 In row-1, the entire image with all the objects is encoded using a global CLIP feature.
 In contrast, row-3 applies a finer-grained approach by masking and encoding each object individually, followed by feature concatenation.
-This object-wise conditions leads to a 1% improvement in object CLIP similarity, indicating better control over individual object appearances.
-Finally, row-4 of all three camera views demonstrate that combining background inpainting with object-wise CLIP encoding achieves the best performance in both foreground and background appearance control.
+These object-wise conditions lead to a 1% improvement in object CLIP similarity, indicating better control over individual object appearances.
+Finally, row-4 of all three camera views demonstrates that combining background inpainting with object-wise CLIP encoding achieves the best performance in both foreground and background appearance control.
 This setup yields the highest CLIP similarity scores for both regions, confirming the effectiveness of jointly modeling structured and unstructured visual components.
-Visualizations.
-We demonstrate that RoboTransfer enables controllable manipulation of both foreground and background appearance while preserving geometric and multi-view consistency across camera views.
-As shown in Figure 4, RoboTransfer successfully modifies the background appearance while retaining the foreground object’s texture and scene geometry.
-Figure 5 illustrates that the model can alter the foreground appearance without affecting the background or spatial layout.
-Finally, Figure 6 demonstrates the model’s ability to jointly edit both foreground and background elements.
-All visualizations confirm that RoboTransfer maintains geometric and multi-view consistency throughout the generated scenes.
+4.2 Synthesis Qualitative Result
+Real-to-Real Synthesis.
+We demonstrate that RoboTransfer enables precise, controllable manipulation of scene elements while preserving global geometric and multi-view consistency.
+As illustrated in Figure 4, our method allows for flexible editing of background attributes (e.g., texture, color) while keeping the foreground and scene geometry intact.
+Similarly, Figure 5 show that foreground appearance can be altered independently of the background.
+Figure 6 further highlights the model’s ability to jointly edit both components.
+By generating diverse variations from structured inputs, this real-to-real framework effectively enriches training data, leading to improved policy generalization for downstream tasks.
+Sim-to-Real Synthesis In simulation, per-view geometry conditions can be obtained at no additional cost.
+As shown in Figure 7, RoboTransfer generates photorealistic videos from simulated geometry inputs, including out-of-distribution cases.
+This Sim-to-Real paradigm reduces reliance on real-world geometry data, thereby better supporting downstream robotic learning tasks.
+Diverse Scene Synthesis For fixed-arm tabletop tasks, backgrounds are simple, while mobile manipulation involves complex geometries.
+With more training data from [8], Figure 8 shows that RoboTransfer can generate richer, more diverse scenes in complex settings, enhancing the variety and realism of synthetic data for robotic learning.
+4.3 Qualitative Comparison
+Figure 9 compares robot data generation methods.
+Cosmos-Transfer [2] performs well under fixed camera views but degrades notably for dynamic viewpoints.
+In contrast, RoboTransfer maintains strong multi-view consistency, producing realistic and coherent novel-view synthesis.
+RoboEngine [55], based on image inpainting, suffers from noise and jitter, lacks temporal consistency, and cannot precisely control scene backgrounds or objects.
 […]
-Figure 4: Visualizations of RoboTransfer with different background reference images.
+Figure 8: Visualizations of RoboTransfer for diversity scene generation results.
 […]
-Figure 5: Visualizations of RoboTransfer with different object reference images.
-[…]
-Figure 6: Visualizations of RoboTransfer with different background and object reference images.
-4.2 Real Robot Experiments
-To evaluate RoboTransfer, we conducted experiments on a real-world robotic platform, assessing how synthetic data impacts visual policy generalization across varied environments.
+Figure 9: Comparison with Cosmos-Transfer and RoboEngine generation results.
+4.4 Comparisons with External Methods
+To the best of our knowledge, RoboTransfer is the first method to enable multi-view consistent data synthesis for robotic manipulation.
+As illustrated in Figure 9, prior approaches suffer from significant multi-view inconsistencies and offer limited background control.
+Furthermore, standard domain randomization techniques fail to model localized or structurally meaningful variations.
+In contrast, RoboTransfer leverages a multi-view diffusion model to generate geometry-consistent, high-fidelity variations in both object appearance and background, thereby supporting robust policy learning.
+5 Real Robot Experiments
+To evaluate the effectiveness of RoboTransfer, we conducted experiments on a real-world robot to assess how synthetic data impacts visual policy generalization.
+Table 4: Effectiveness of Synthetic Data Augmentation. Performance comparison across different data augmentation strategies. Success Rate (%) and Score are reported.
+Data Composition Spoon Pick&Place Towel Folding
+Diff-Obj Diff-All Diff-Obj Diff-All
+SR Score SR Score SR Score SR Score
+Real only 33.3% 2.67 13.3% 1.56 16.7% 1.81 12% 1.08
+Domain Random Aug 44.4% 2.89 11.1% 1.58 16.7% 1.92 12% 1.12
+Real + Obj Aug 44.4% 3.00 22.2% 2.04 16.7% 2.00 12% 1.24
+Real + Obj&Bg Aug 66.7% 3.56 46.7% 2.98 50.0% 2.60 28% 1.92
+5.1 Experimental Platform and Implementation
 Experimental Platform and Setup Details.
-Our experiments were conducted on the Agilex Cobot Magic platform, equipped with two PIPER robotic arms and three Intel RealSense D435i cameras—two mounted on the wrists and one positioned overhead.
-We used only RGB data for policy learning, creating a challenging visual understanding problem while reflecting realistic deployment constraints.
-Benchmark Task and Evaluation Methodology.
-We selected a dual-arm manipulation sequence with four phases: (i) left arm grasps a spoon; (ii) spoon is placed on a tray; (iii) right arm grasps the spoon; (iv) spoon is placed in a cup.
-We evaluated across three test conditions: identical objects/background (Same), novel object appearances (Diff-Obj), and variations in both objects and environment (Diff-All).
-Testing protocols are detailed in the supplementary material.
-We introduced a Stage Score (0-4) corresponding to successfully completed phases, providing finer-grained insights than binary success/failure metrics.
-This reveals how far policies progress through the task sequence, particularly valuable when analyzing generalization.
+Our experiments were conducted on the Agilex Cobot Magic platform, which is equipped with two PIPER robotic arms and three Intel RealSense D435i cameras, two mounted on the wrists and one positioned overhead.
+Only RGB data was used for policy learning, presenting a challenging visual understanding problem.
 Implementation Details.
@@
 For each task, we collected 100 expert demonstrations via ALOHA [61] teleoperation.
-Our methodology follows a pretrain+finetune paradigm.
-We trained baseline policies on real data for 100k steps, then fine-tuned with synthetic data for 50k additional steps (learning rate reduced to 1/10).
-For the grasp spoon task, training on real data required 8 NVIDIA H20 GPUs for approximately 24 hours, while the synthetic data fine-tuning phase utilized the same hardware configuration for about 12 hours.
-Additional details of the training protocol are given in the supplementary material.
-Effectiveness of Synthetic Data Augmentation.
-Table 3: Performance comparison across different test conditions and models.
-Success Rate (%) and Stage Score (0-4) are reported for each configuration.
-The best results for each test conditions are highlighted in bold.
-Model Same Diff-Obj Diff-All
-Success (%) Score Success (%) Score Success (%) Score
-Baseline (Real only) 80.0 3.6 33.3 2.7 13.3 1.6
-+ Obj Augmentation 44.4 3.0 22.2 2.0
-+ Obj+Bg Augmentation 46.7 3.0
-Table 3 presents the comparative results.
-Results show significant performance degradation when baseline policies face novel visual contexts, with success rates dropping from 80.0% to 13.3%.
-However, policies fine-tuned with our generated data show considerable resilience.
-In Diff-Obj conditions, success rates improved by $33.3\%$ (relative), while in Diff-All scenarios, we observed a $251\%$ relative improvement.
-Stage Score results further confirm that augmented models make more progress through task phases, indicating more robust manipulation understanding even when tasks aren’t fully completed.
+To demonstrate generalizability, the scenes and tasks differ from those in the RoboTransfer training data (see Sec.
+4 ).
+More details are provided in the appendix.
+5.2 Effectiveness of Synthetic Data
+Benchmark Task and Evaluation Methodology.
+We evaluate synthetic data on two long-horizon dual-arm manipulation tasks: spoon pick-and-place and towel folding.
+Spoon pick-and-place consists of four phases: the left arm grasps a spoon, places it on a tray, the right arm grasps it, and places it in a cup.
+Towel folding includes three stages: grasping the bottom corners, lifting and folding upward, and folding the right side to the left.
+The towel task provides a rigorous test of physical plausibility and temporal coherence in complex manipulations.
+We introduce a Stage Score to measure phase completion, offering finer-grained insight than binary success metrics and enabling evaluation of policy progression and generalization.
+Experiments were conducted under two conditions: novel objects (Diff-Obj), and variations in both objects and environment (Diff-All).
 Effect of Synthetic Data Proportions.
-We investigated how different proportions of synthetic data affect performance by evaluating synthetic ratios of 25%, 50%, 75%, and 100%.
-Figure 7 shows these results for the Diff-All conditions.
-Performance follows a non-linear relationship with synthetic data proportion, peaking at 50% and showing diminishing returns beyond.
-Success rates increased from 13.3% with no synthetic data to 46.7% at 50% synthetic ratio—a 3.5x improvement.
-Stage scores similarly peaked at 3.0 with 50% mixture versus 1.6 without synthetic data.
-Notably, using exclusively synthetic data (100%) yielded 40.0% success rate and 2.9 stage score—improvements over purely real data but below the balanced 50% mixture.
-This confirms complementary contributions: synthetic data provides diversity while real data contributes physical fidelity, with optimal interaction at equal proportions.
-RoboTransfer preserves geometric consistency while enabling systematic appearance variation, addressing a fundamental challenge in robot learning.
-Additional details on model configuration and extended evaluations are in the supplementary material.
-Figure 7: Performance metrics across synthetic data mixing ratios.
-Left: Success rate vs.
-ratio.
-Right: Average stage score vs.
-ratio.
-Note that 0% represents our baseline model trained only on real data.
-Both metrics peak at 50% synthetic data, suggesting an optimal balance between synthetic diversity and real-world physical fidelity.
+To determine the optimal data mixture, we evaluated policy performance on the spoon pick-and-place task under the Diff-All setting, varying the synthetic data proportion from 0% to 100%.
+As shown in Figure 10, both metrics peak at a 50/50 ratio.
+With this balanced distribution, the success rate increases significantly from 13.3% to 46.7%, and the stage score nearly doubles from 1.6 to 3.0.
+Further increasing the synthetic proportion leads to performance degradation; although geometrically consistent, synthetic data may lack fine-grained physical fidelity (e.g., contact dynamics or material properties), which becomes detrimental at excessive ratios.
+Notably, a policy trained exclusively on synthetic data still achieves a 40.0% success rate, surpassing the real-only baseline.
+This highlights the complementary roles of the two sources: synthetic data provides visual diversity, while real data grounds the policy in real-world physical dynamics.
+Consequently, we adopt the 50/50 ratio for the analysis in the Effectiveness of Synthetic Data section.
+Figure 10: Performance across synthetic data mixing ratios. Note 0% represents the baseline model trained only on real data.
+Effectiveness of Synthetic Data Analysis.
+As shown in Table 4, our synthetic data augmentation significantly enhances policy robustness where the baseline fails.
+For the Spoon Pick and Place task under the most difficult Diff-All condition, our approach achieves a 251% relative improvement in success rate (from 13.3% to 46.7%).
+The results also reveal a clear cumulative benefit of our augmentation strategy.
+In the Towel Folding Diff-Obj setting, for instance, the Stage Score increases from 1.81 (real only) to 2.0 with object augmentation, and further to 2.6 after adding background augmentation.
+This confirms that augmenting both object and background appearance is key to learning policies that can handle significant visual domain shifts.
+While domain randomization handles object color changes, it cannot model localized or structurally meaningful variations.
 6 Conclusion
-In this work, we proposed RoboTransfer, a novel diffusion-based data synthesis framework in robotics, which integrates multi-view geometry while offering explicit control over background and object attributes.
-We also introduced our dataset processing pipeline, which provides triplets incorporating both global geometry and appearance control conditions.
-The experiments and evaluations show that our RoboTransfer can produce multi-view consistency data for training robotic manipulation policies.
-Therefore, the generalization of visuomotor policy gets significantly improved.
-Limitation.
-Although the method is general and flexible, several limitations remain that we aim to address in future work.
-First, our approach currently focuses on augmenting the appearance of objects and scenes, rather than generating new motions.
-Moving forward, we plan to integrate RoboTransfer with additional simulators to broaden its applicability.
-Furthermore, we aim to scale up both the size and diversity of the training data, which will enhance the variety and robustness of the generated outputs.
-Appendix
+In this work, we proposed RoboTransfer, a diffusion-based data synthesis framework for robotics that integrates multi-view geometry while providing explicit control over background and object attributes.
+We also introduced a dataset construction pipeline that generates high-quality triplets incorporating both global geometry and appearance conditions.
+Experimental results demonstrate that RoboTransfer produces multi-view consistent data, significantly improving the generalization of visuomotor policies for robotic manipulation.
 […]
-1
-Appendix A Dataset Construction Details
-For the training of RoboTransfer, we collected data and designed a dedicated dataset construction pipeline (see Figure 3). In this section, we provide additional implementation details.
-A.1 Data Collection
-Robot Platform: We construct a large-scale robotic demonstration dataset using the Agilex Cobot Magic platform, adhering to standardized data collection protocols established in prior work [44, 39].
-The robot is equipped with two PIPER robotic arms.
-As illustrated in Figure 8, each demonstration includes synchronized RGB-D streams captured from three Intel RealSense D435i cameras: one mounted at two hand-eye viewpoint and one positioned overhead.
-Dataset Design: To capture a broad spectrum of manipulation scenarios, we systematically vary tasks, backgrounds, and manipulable objects.
-Specifically, we design twelve distinct manipulation tasks (visualized in Figure 9), each representing diverse interaction patterns.
-For each task, we collect 100 demonstration segments across 10 unique object configurations, resulting in a total of 1,000 samples per task.
-The dataset incorporates a variety of backgrounds, including textured tabletops and cluttered surfaces, to reflect realistic operating conditions.
-Manipulable objects are diversified in terms of shape, size, and material, enhancing both the visual richness and physical complexity of the dataset.
-This comprehensive design ensures a high degree of diversity and realism, which is crucial for training robust and generalizable visual policies.
+Supplementary Material
+7 RoboTransfer Implementation Details
+7.1 Training Details
+Training Dataset.
+To construct our training dataset, we leverage the open-source Cobot Magic platform to collect a large-scale video corpus of dual-arm robot executions(see Sec.
+8.1 for details).
+The raw videos are segmented into 10Hz clips of 30 frames each, resulting in approximately 24k clips for training.
+Additionally, we curate a set of 1.6k 10Hz 30-frame clips from the collected dataset for video synthesis quality evaluation.
+All videos are annotated with conditions to facilitate both training and assessment.
+Training Details.
+RoboTransfer is fine-tuned from the pre-trained Stable Video Diffusion [5] model.
+During training, videos from each camera view are resized to a resolution of $640\times 384$.
+We adopt the AdamW[35] optimizer with a learning rate of $3\times 10^{-5}$ and a global batch size of 8, training for a total of 70K steps.
+During inference, we use the EDM scheduler [24] to perform 30 denoising steps and apply classifier-free guidance.
+7.2 Modeling and Conditions Injection Details
+Object Condition.
+We resize each object image to $224\times 224$ and pass it through CLIP’s image encoder to obtain a single global feature vector per object.
+These embeddings are then concatenated and fed into the diffusion model.
+This design allows for finer-grained control, enabling individual manipulation of each object.
+Multi-View Consistency Modeling.
+Previous methods typically introduce a cross-view module to enhance consistency between views.
+In contrast, RoboTransfer simply concatenates multi-view images, integrating inter-view consistency into global spatial consistency, thereby improving multi-view video modeling.
+Moreover, this design allows direct loading of pre-trained single-view video generation model weights, without requiring significant modifications to the model architecture.
+8 RoboTransfer Dataset Construction Details
+For training RoboTransfer, we collected a dedicated dataset and designed a construction pipeline (Figure 3 in the main paper). Here, we provide additional implementation details.
+8.1 Data Collection
+Robot Platform.
+We built a large-scale robotic demonstration dataset using the Agilex Cobot Magic platform, following standardized protocols [37, 33].
+Each demonstration includes synchronized RGB-D streams from three Intel RealSense D435i cameras: two hand-eye views and one overhead view (Figure 12).
+Dataset Design.
+To capture diverse manipulation scenarios, we designed twelve distinct tasks (Figure 12) with variations in objects, backgrounds, and interactions.
+For each task, 100 demonstration segments were collected across 10 unique object configurations, totaling 1,000 samples per task.
+Backgrounds range from textured tabletops to cluttered surfaces, and objects vary in shape, size, and material.
+This ensures high diversity and realism for robust visual policy training.
+To further enhance background diversity, we incorporate the AgiBot-World dataset [8].
 […]
-Figure 8: Visualizations of robot platform.
+Figure 11: Robot platform visualizations.
 […]
-Figure 9: Visualizations of the dataset.
-A.2 Geometry conditions
-Depth conditions: To enforce both spatial and temporal consistency, we incorporate geometry-conditioned supervision into our dataset construction pipeline.
-During data collection, we obtain synchronized RGB-D streams, including multi-view depth maps.
-However, depth maps from the RGB-D sensors are often noisy and incomplete, particularly in regions with low reflectivity (e.g., black surfaces), leading to domain gaps across samples.
-In contrast, depth maps rendered from simulators are typically complete and noise-free.
-While we explored existing depth completion methods, we found that many suffer from limited temporal consistency and poor generalization across diverse scenes.
-To address these limitations, we adopt the state-of-the-art Video Depth Anything (VDA) model [9], which produces temporally coherent and spatially complete depth maps.
-Despite the high-quality temporal consistency of depth estimates from Video Depth Anything (VDA), these predictions lack a globally consistent scale across multiple views.
-To address this, we align the VDA-predicted depth maps with the RGB-D sensor measurements via a robust least-squares scale fitting procedure.
-While standard least-squares regression offers a closed-form solution for affine scale alignment, it is susceptible to errors introduced by sensor noise and depth outliers.
-To mitigate this, we introduce a robust multi-frame fitting strategy (Figure 10) that aggregates depth correspondences across time and iteratively filters outliers based on residual error thresholds.
-This iterative refinement ensures stable and accurate recovery of the global scale transformation.
-Our approach effectively preserves the geometric completeness provided by VDA while anchoring depth predictions to the real-world scale captured by the RGB-D sensor.
-The resulting alignment ensures both spatial accuracy across views and temporal smoothness across frames, enabling reliable depth conditions.
-Input:
-Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
-Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
-Output:
-Scaled Depths $\mathbf{D}_{\text{scaled}}\in\mathbb{R}^{B\times H\times W}$,
-1
-2Initialize Mask: $\mathcal{M}\leftarrow(\mathbf{D}_{\text{sensor}}>\epsilon)\land(\mathbf{D}_{\text{pred}}>\epsilon)$ ;
-3
-4for 2 iterations do
-5       $\mathcal{E}\leftarrow|\mathbf{D}_{\text{pred}}-\mathbf{D}_{\text{sensor}}|\odot\mathcal{M}$ ;
-6       $\tau\leftarrow\mathcal{P}_{80}(\mathcal{E}[\mathcal{M}>0])$ ;
-7       $\mathcal{M}\leftarrow(\mathcal{E}<\tau)\odot\mathcal{M}$ ;
-8       $s,b\leftarrow\textbf{Scale Fitting}(\mathbf{D}_{\text{pred}},\mathbf{D}_{\text{sensor}},\mathcal{M})$ ;
-9       $\mathbf{D}_{\text{scaled}}\leftarrow s\cdot\mathbf{D}_{\text{pred}}+b$ ;
-10
-return $\mathbf{D}_{\text{scaled}}$;
-Algorithm 1 Dynamic Mask Refinement
-Input:
-Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
-Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
-Mask $\mathcal{M}\in\{0,1\}^{B\times H\times W}$
-Output: Scale $s\in\mathbb{R}$, Shift $b\in\mathbb{R}$
-1 Flatten and Mask:
-$\mathbf{p}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{pred}}\odot\mathcal{M})\in\mathbb{R}^{N}$
-$\mathbf{s}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{sensor}}\odot\mathcal{M})$
+Figure 12: Dataset collected with the Agilex robot.
+8.2 Geometry Conditions Construction
+Depth Conditions.
+For RGB-D data, raw sensor depth provides accurate metric information but is often noisy or incomplete, particularly on low-reflectivity surfaces.
+Conversely, simulator-rendered depth is perfect but lacks domain realism.
+To bridge this gap, we adopt Video Depth Anything (VAD) [9] to generate temporally coherent and spatially complete depth maps.
+However, VAD predicts relative depth, lacking the global metric scale essential for robotic manipulation.
+We therefore treat the raw sensor depth as a sparse metric anchor and the VAD output as a dense structural prior.
+We align them using a robust multi-frame least-squares fitting strategy (Figure 13).
+Crucially, this process employs an iterative outlier filtering mechanism (Algorithm 1) that allows the model to learn the correct scale $s$ and shift $b$ from reliable sensor pixels while ignoring sensor noise/holes.
+This ensures the final output inherits the metric accuracy of the sensor while retaining the completeness of VAD.
+For datasets without multi-view RGB-D sensors, such as AgiBot-World [8], we estimate metric depth using MoGe [48].
+Algorithm 1 Dynamic Mask Alignment
 […]
-2Construct Global System:
-$\begin{bmatrix}\mathbf{p}^{\top}\mathbf{p}&\mathbf{p}^{\top}\mathbf{1}\\ \mathbf{p}^{\top}\mathbf{1}&\mathbf{1}^{\top}\mathbf{1}\end{bmatrix}\begin{bmatrix}s\\ b\end{bmatrix}=\begin{bmatrix}\mathbf{p}^{\top}\mathbf{s}\\ \mathbf{1}^{\top}\mathbf{s}\end{bmatrix}$
-Closed-Form Solution:
-$s=\frac{(\mathbf{1}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{1}^{\top}\mathbf{s})}{\det}$
-$b=\frac{(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})}{\det}$
-where $\det=(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{1})-(\mathbf{p}^{\top}\mathbf{1})^{2}>\epsilon$
-Algorithm 2 Scale Fitting
-Figure 10: Depth Scale Alignment: Implementation Details
-Normal conditions: For high-precision surface reconstruction, surface normals are critical and can provide surface details rather than only depth.
-At the same time, normal does not have scale consistency issues.
-We compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS [24].
-A.3 Appearance conditions
-Object Mask: For the manipulation task data, we first select keyframes that effectively capture both the objects and the background conditions of the scene.
-We then extract all operable objects using a structured prompt template (see Figure 11), which describes each object in terms of its color, material, shape, and spatial position.
-These object descriptions are fed into Grounding DINO[38] to obtain bounding boxes for object detection.
-The resulting bounding boxes are subsequently used as prompts for SAM2[49] to generate the corresponding object masks.
-*System Instruction*
-As a visual descriptor, articulate objects with this flow:
-Color -> Material -> Shape/Texture -> Spatial Position
-*Template Format*
-"A [color] [material] [object], [shape] with [texture], located [region](x,y)"
-*Demonstration Examples*
-- A red plastic ball, spherical with a smooth glossy surface, at the center (250,300)
-- A brown wooden chair, angular with a matte wood grain, in top-left (100,50)
-- A silver-blue metallic arm, mechanical with a brushed texture, on the right (600,200)
+[…]
+3:Initialize mask:
+4:   $\mathcal{M}\leftarrow(\mathbf{D}_{\text{sensor}}>\epsilon)\land(\mathbf{D}_{\text{pred}}>\epsilon)$
+5:for $i=1$ to 22 do
+6:  Scale Fitting:
+7:  $s,b\leftarrow\text{ScaleFitting}(\mathbf{D}_{\text{pred}},\mathbf{D}_{\text{sensor}},\mathcal{M})$
+8:  $\mathbf{D}_{\text{metric}}\leftarrow s\cdot\mathbf{D}_{\text{pred}}+b$
+9:  Mask Update:
+10:  $\mathcal{E}\leftarrow|\mathbf{D}_{\text{pred}}-\mathbf{D}_{\text{sensor}}|\odot\mathcal{M}$
+11:  $\tau\leftarrow\text{Percentile}_{80}(\mathcal{E}[\mathcal{M}>0])$
+12:  $\mathcal{M}\leftarrow(\mathcal{E}<\tau)\odot\mathcal{M}$
+13:end for
+14:return $\mathbf{D}_{\text{metric}}$
+Algorithm 2 Scale Fitting (Least Squares)
+[…]
+[…]
+3:Extract valid pixels:
+4:   $\mathbf{p}=\mathbf{D}_{\text{pred}}[\mathcal{M}],\quad\mathbf{s}=\mathbf{D}_{\text{sensor}}[\mathcal{M}]$
+5:Objective:
+6:   $\min_{s,b}\|s\mathbf{p}+b\mathbf{1}-\mathbf{s}\|^{2}$
+7:Closed-form Solution:
+8:    $\begin{bmatrix}s\\ b\end{bmatrix}=\frac{1}{\Delta}\begin{bmatrix}N(\mathbf{p}^{\top}\mathbf{s})-(\mathbf{1}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})\\ (\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{s})(\mathbf{1}^{\top}\mathbf{p})\end{bmatrix}$
+9:   where $\Delta=N(\mathbf{p}^{\top}\mathbf{p})-(\mathbf{1}^{\top}\mathbf{p})^{2}$
+10:return $s,b$
+Figure 13: Depth Scale Alignment Algorithms.
+We employ an iterative robust least-squares fitting strategy.
+Algorithm 1 aligns relative depth predictions to sparse metric sensor data by iteratively filtering high-error regions (outliers).
+Algorithm 2 provides the analytical closed-form solution for the linear alignment parameters $s$ (scale) and $b$ (shift).
+Normal Conditions Surface normals capture fine geometric details and are scale-invariant.
+We compute per-frame normals using LOTUS [20].
+For datasets lacking multi-view RGB-D sensors, MoGe [48] simultaneously estimates depth and normals, streamlining prelabeling.
+8.3 Appearance Conditions Construction
+Keyframe Selection.
+Keyframes serve as the reference for object and background appearance.
+For standard tabletop tasks, we adopt a temporal strategy: the initial frame (fully populated) defines the object condition, while the final frame (post-manipulation) serves as the background.
+For complex scenarios, we employ an automated content-aware pipeline using VLM descriptors and Grounding DINO.
+This pipeline selects the frame maximizing the object’s visible pixel area for the object condition, and conversely, the frame with minimal object presence for the background condition.
+Object Mask Generation.
+Object descriptions, generated via structured prompts (Figure 14) specifying color, material, shape, and spatial position, are fed into Grounding DINO [34] to generate bounding boxes, which SAM2 [41] converts into object masks.
+System Prompt for Object Descriptor: You are an industrial robotic vision system with 100% detection guarantee.
+Perform comprehensive scene analysis to identify ALL movable objects, including occluded items.
+Strictly enforce size constraints and occlusion handling.
+Your goal is to complete a list of all visible objects on the table and process the description.
+The output Template Format: A [color] [object], [shape], located [region](x,y) Examples: A red ball, spherical, at the center (250,300); A brown chair, angular, in the top-left (100,50); A silver-blue arm, mechanical, on the right (600,200);
 Figure 14: Visual Description Prompt Template Architecture.
-Objects and Background conditions:To extract the object conditions, we first segment image patches corresponding to individual objects from the full image.
-Each patch is resized to 384×384 and passed through the CLIP model [47] to obtain RGB embeddings.
-For the background conditions, we mask out all detected objects from the full image using their corresponding masks.
-The masked image is then fed into an inpainting model to reconstruct the background with the objects removed.
-Appendix B Visual Policy Implementation Details
+Objects and Background Conditions.
+Individual object patches are resized to $224\times 224$ and passed through CLIP [40] to obtain embeddings.
+Background conditions are generated by masking out all detected objects and applying inpainting to reconstruct the object-free scene.
+9 Robot Policy Model Implementation Details
 To validate the effectiveness of the data generated by RoboTransfer, we train a visual policy using the procedures detailed below for both training and deployment.
-B.1 Data Collection and Preprocessing
-To ensure the validity of our evaluation, we excluded the dataset described in Section A from the real-robot experiments. The data preparation proceeded as follows:
-Real Expert Data: We collected 100 sets of expert demonstration data for each manipulation task using the ALOHA teleoperation system.
-The primary observations were RGB images (1280x720 resolution, subsequently downscaled to 640$\times$360 to optimize training), with camera streams captured at approximately 30Hz and sampled at 10Hz for policy training.
-Auxiliary robot state information (e.g., joint positions, end-effector poses) was originally recorded at 200Hz and downsampled to 50Hz for model use.
-Synthetic Data Generation: To further enhance model performance and generalization, we generated synthetic data based on the collected real data, specifically targeting variations in foreground objects and background scenes, for use in the subsequent fine-tuning phase.
-B.2 Training Pipeline and Core Parameters
-Our training pipeline is based on the ACT (Action Chunking Transformer) architecture, which processes visual input from three distinct perspectives (two wrist-mounted cameras and one fixed overhead camera).
-At each training timestep, the model receives a single RGB frame from each view.
-Training Objective and Two-Stage Training Strategy: The policy model is designed to predict a sequence of the next 100 robot states, corresponding to a 2-second action horizon given the 50Hz robot state frequency.
-We adopted a "pre-train-then-finetune" paradigm:
-Pretraining on Real Data:The model was first pretrained for 100k steps using the collected real expert demonstration data.
+9.1 Data Collection and Preprocessing
+To ensure a fair evaluation, we excluded the RoboTransfer training dataset (Sec. 8.1) from the real-robot experiments. Data preparation was conducted as follows:
+Real Expert Data We collected 100 expert demonstration sets per manipulation task using the ALOHA teleoperation system.
+Observations included RGB images at 1280×720 resolution, downscaled to 640×360 for training efficiency, captured at 30Hz, and sampled at 10Hz.
+Auxiliary robot state information, including joint positions and end-effector poses, was recorded at 200Hz and downsampled to 50Hz for policy training.
+Synthetic Data Generation for Policy Fine-tuning.
+To improve generalization, we generated synthetic videos based on the real demonstrations, introducing variations in foreground objects and background scenes.
+The pretrained diffusion model conditioned each synthetic video on: (1) per-frame 3D geometry inputs (depth and normal maps) from real demonstrations, and (2) a reference image containing novel objects and backgrounds from a held-out set independent of the policy training data.
+9.2 Training Pipeline and Core Parameters
+Our training pipeline uses the ACT (Action Chunking Transformer [60]) architecture, processing visual input from three cameras (two wrist-mounted, one overhead).
+At each timestep, the model receives one RGB frame per view.
+Training Objective and Strategy.
+The policy predicts the next 100 robot states (2s horizon at 50Hz).
+We adopt a pretrain-then-finetune strategy: 1) Pretraining on Real Data: The model was first pretrained for 100k steps using the collected real expert demonstration data.
 During this phase, the batch size was set to 512, and the learning rate was $1\times 10^{-4}$.
-Finetuning with Synthetic Data: After pretraining, the aforementioned synthetic data was introduced to fine-tune the model for an additional 50k steps.
+2) Finetuning with Synthetic Data: After pretraining, the synthetic data was introduced to fine-tune the model for an additional 50k steps.
 The learning rate for this phase was reduced to $1\times 10^{-5}$.
 All training was performed on a cluster equipped with 8 NVIDIA H20 GPUs. The pretraining phase took approximately 24 hours, and the fine-tuning phase required about 12 hours.
-B.3 Real-Robot Deployment and Evaluation
-The trained policies were deployed and evaluated in the following environment:
-Deployment Platform: Experiments were conducted on the Agilex Cobot Magic platform8, the same platform used for initial data collection, ensuring consistency between training and testing environments.
-Inference Procedure: During deployment, the policy operated in a synchronous execution mode.
-Specifically, the system waits for the robot to fully execute the action sequence (comprising 100 actions) generated from the model’s previous inference step before capturing new observations and proceeding to the next model inference and subsequent execution.
-At each decision point, the model receives as input a single RGB frame from each of its multiple camera views and the current robot state and outputs the sequence of 100 actions.
-The latency for this single model inference process is approximately 10ms.
-These actions are then executed on the robot at a 50Hz control frequency, consistent with the robot state prediction rate during training.
-Appendix C More Qualitative Results
-C.1 Real-to-Real Transfer
-The proposed RoboTransfer offers strong controllability over both foreground and background generation.
-Given the same structured input, the model allows for flexible editing of background attributes such as texture and color, as shown in Figure 12.
-Meanwhile, as illustrated in Figure 13, the appearance of foreground objects, including their color, can also be effectively modified.
-This real-to-real generation framework enriches the diversity of training data for downstream tasks, contributing to better generalization of the policy model.
-[…]
-Figure 12: Visualizations of RoboTransfer with different background reference images.
-[…]
-Figure 13: Visualizations of RoboTransfer with different object reference images.
-C.2 Sim-to-Real Transfer
-As shown in Figure 14, RoboTransfer generates photorealistic videos from simulated structural inputs, including out-of-distribution cases.
-This Sim-to-Real paradigm reduces dependency on structured information from real-world datasets, thereby better supporting downstream robotic learning tasks.
-[…]
-Figure 14: Visualizations of Sim-to-Real generation results.
-C.3 Comparison
-Figure 15 illustrates that Cosmos Transfer [2] is mainly effective for robot scene generation under fixed camera settings (central view), while its quality drops notably for dynamic camera viewpoints (side views).
-Compared to Cosmos Transfer [2], RoboTransfer demonstrates significantly better multi-view consistency, enabling more realistic and coherent novel-view synthesis across varying camera poses.
-[…]
-Figure 15: Comparison with Cosmos generation results.
+9.3 Real-Robot Deployment and Evaluation
+Deployment Platform.
+Policies were evaluated on the Agilex Cobot Magic platform (Figure 12), the same system used for data collection to ensure consistency between training and evaluation environments.
+Inference Procedure During deployment, the policy operates synchronously: the robot executes the full action sequence (100 actions) generated from the previous inference step before capturing new observations.
+At each decision point, the model receives one RGB frame per camera view along with the current robot state and outputs the next 100 actions.
+Inference latency is  10ms, and actions are executed at 50Hz, matching the training robot state frequency.
+10 Discussion
+10.1 Analysis on Geometry-Appearance Decoupling
+In our framework, reference images serve primarily as appearance conditions, capturing visual properties like texture and color, rather than imposing strict geometric constraints.
+Global 3D geometry conditions ensure multi-view and temporal consistency across frames.
+1.
+Geometry conditions.
+Multi-view depth and normal maps encode precise 3D structure and scene layout, even under dynamic wrist-mounted camera motion, defining where and how content should be rendered.
+2.
+Appearance transfer.
+The reference image provides the target visual style (e.g., textures or background appearance) that the diffusion model transfers onto the global geometry.
+The reference does not need to be aligned with each frame; the model synthesizes a consistent appearance across views and time by using geometry as an anchor.
+Qualitative results show visually coherent, realistic videos in wrist-view settings.
+Real-to-Real experiments confirm effective appearance transfer to dynamic robotic scenes, while Sim-to-Real experiments demonstrate that geometry from simulation can be faithfully combined with real-world appearance references.
+By decoupling geometry and appearance, our method improves robustness and generalization, requiring only diverse reference images rather than per-scene 3D scans or carefully aligned inputs, making it scalable for real-world deployment.
+10.2 Scalability of 3D Data Acquisition
+Globally consistent 3D conditions are essential for wrist-view camera motion.
+While acquiring dense, high-quality 3D data can be resource-intensive, our framework leverages automated, cost-efficient pipelines:
+1.
+Real-robot data: RGB-D cameras capture metric depth alongside RGB images.
+For both RGB-D and RGB-only images, depth and normal maps are robustly generated using automated pipelines.
+The code will be released to ensure reproducibility.
+2. Simulation data: Per-view 3D geometry conditions (depth and normals) are trivially available at no extra cost, making simulation a scalable source of 3D-aware inputs.
+By combining real and simulated pipelines, our framework reduces the effective cost of 3D data acquisition, enabling high diversity and fine-grained controllability in synthetic video generation for robotic policy learning.
+10.3 Rationale for Reference Selection
+RoboTransfer is designed for scalable, fully automated training.
+Capturing a clean background for every demonstration would be prohibitively expensive.
+Instead, background imagery is extracted from existing videos, preserving fidelity without extra data collection.
+In inference, a clean background frame can be captured if desired, keeping the process simple and efficient.
+This design balances automation, scalability, and flexibility across training and deployment.